### v2

# Quantum Hermitian conjugate and encoding unnormalized matrices

Abstract
We further develop the family of matrix-manipulation algorithms based on the encoding the matrix elements into the probability amplitudes of the pure superposition state of a certain quantum system. We introduce two extensions to these algorithms which allow (i) to perform Hermitian conjugation of matrices under consideration and (ii) to weaken the restriction to the absolute values of matrix elements unavoidably imposed by the normalization condition for a pure quantum state. Both these extensions are applied to the matrix multiplication algorithm. Controlled measurement of ancilla state is implemented to avoid the problem of small success probability in the measurement process.
I Introduction
Quantum algorithm is a rapidly developing area of quantum informatics whose progress is stimulated by prospective enormous advantages of quantum algorithms over their classical counterparts. We have to acknowledge such famous algorithms as the Deutsch algorithm Deu demonstrating effect of quantum parallelism in operation with Boolean functions, Shor’s factoring algorithm Sho1 ; Sho2 utilising the quantum Fourier transform QFT1 ; QFT2 ; QFT3 , Grover’s search algorithm Gro , phase-estimation algorithm for the Hermitian QFT3 and non-Hermitian Wang matrices .
We consider a particular application area of quantum algorithms which covers matrix algebraic operations, such as multiplication, addition, determinant calculation, matrix inversion and its application to linear system solver. During recent years, there are many quantum algorithms solving various aspects of linear algebra. Among them are HHL-algorithm for solving systems of linear equations HHL which was applied, in particular, to Gaussian process regression ZhaoZ , algorithm for solving linear differential equations Berry , algorithm for evaluation of matrix functions Tak .
A possible approach to quantum algorithms for matrix manipulations is proposed in ZZRF . This approach refers to matrices as operators to be applied to some vector. Therefore, it is essentially in that approach that arbitrary nonunitary and non-Hermitian matrix must be transferred to the unitary one. Thus, first one has to replace a given matrix $A$ (rectangular in general) with the Hermitian one and then transfer it to the unitary operator by exponentiating which can be approximated via Trotterization Trotter ; Suzuki ; T ; AT and Baker-Champbell-Hausdorff method BC . To implement such matrix transformations effectively, the set of special matrices is proposed to be prepared in advance. The phase estimation subroutine QFT3 is included into that algorithm.
A special block-encoding model GSLW is another approach for implementing matrix algebra. Different type of block-encoding was used in the algorithm in Ref.DFZ_2020 to embed the inverse of the matrix into the unitary transformation. In LWWZ , matrix multiplication was performed trough the binary encoding of matrix elements into the pure states of quantum systems with subsequent binary multiplication via Toffoli gates. Matrix multiplication over some rings, in particular, multiplication of Boolean matrices, was studied in KN .
Some algorithms for matrix inversion have been also developed. As the first one we call the mentioned above HHL algorithm HHL where the inversion was used for solving a linear system. However, the problem of inversion of matrix eigenvalues has not been resolved in frames of quantum algorithm therein. The matrix inversion algorithm based on the singular value decomposition is presented in MRTC ; NJ . In both paper the authors apply the function evaluation problem GSLW to approximate the inverse of singular values by odd-polynomial. This approach requires introducing a special scale parameter depending on the interval where all nonzero singular values are distributed.
Recently the alternative matrix-manipulation algorithms based on the matrix encoding into the complex probability amplitudes of a pure superposition state of a certain subsystem have been developed QZKW_arxive2022 ; ZQKW_2024 ; ZBQKW_arXive2024 . The principal difference of such approach is that a matrix is considered as a part of the quantum state of a system rather then a unitary operator applied to this system. Therefore the elements of the resulting matrix (obtained after addition or multiplication of other matrices, or after matrix inversion) can be measured up to some constant factor. Although those algorithms deal with complex matrices in general, some operations (including Hermitian conjugation) can not be realized in frames of those settings and therefore certain constraints are to be imposed on the variety of matrix manipulations. To remove this disadvantage and construct more flexible algorithms for matrix algebra we propose to separate the real and imaginary parts of a matrix by encoding them in different probability amplitudes of a pure state. This separation can be done by involving additional one-qubit subsystem whose state serves to label the real and imaginary parts of a matrix elements. As a consequence, all probability amplitudes are real in such representation.
Another issue addressed in this paper is weakening the normalization constraint for the elements of the considered matrices. We recall that such constraint is imposed by the normalization condition for a pure quantum state. For that purpose we introduce an additional probability amplitude into the superposition state encoding considered matrix. This can be done involving one more auxiliary one-qubit subsystem.
Of course, both above extensions require modifications of the algorithms proposed in ZQKW_2024 ; ZBQKW_arXive2024 . Below we present the detailed analysis of the appropriate modifications introduced into the multiplication algorithm (see appendix in ZBQKW_arXive2024 ) and demonstrate manipulations with the input matrices involving Hermitian conjugation. Similar modifications caused by the above extensions can be implemented in other algorithms such as addition, determinant calculation, matrix inversion, linear system solving. Such modifications will be considered in a different paper.
The paper is organized as follows. In Sec.II we introduced two pointed above extensions to the algorithms based on the matrix encoding into a pure state. In Sec.III, we implement these extensions into the matrix multiplication algorithm, demonstrate possibility of manipulating input matrices using subroutine of Hermitian conjugation and show how to measure the output normalization constant that is loosed because of the controlled measurement FZQW_arxive2025 of the ancilla state at the last stage of the algorithm. Conclusions are given inSec.IV.
II Extensions of matrix encoding technique
The matrix encoding implemented in the algorithms for matrix manipulations proposed in Ref.ZQKW_2024 (see also appendix in (ZBQKW_arXive2024 ) )uses two quantum subsystems $R$ and $C$ whose basis states label the rows and columns associated with each element $a_{jk}$ (complex in general) of the $N\times N$ matrix $A=\{a_{jk}:j,k=0,\dots,N-1\}$. Thus, the matrix $A$ is encoded as follows:
$|\Psi_{A}\rangle$ $=$ $\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}a_{jk}|j\rangle_{R}|k\rangle_{C}$ (1)
with the following normalization condition for the pure state of a quantum system:
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}=1.$ (2)
We introduce two extensions to the algorithms of matrix manipulations developed in ZQKW_2024 . First of them is aimed on effective operations with complex matrices, Sec.II.1. The second one allows to reduce the constraint (2) imposed on the matrix elements, Sec.II.2. We deal with the square $N\times N$ matrices for convenience. If the original matrix is rectangular we turn it to the square matrix by adding the appropriate number of either zero rows or zero columns. We also require $N=2^{n}$, $n$ is a positive integer.
II.1 Hermitian conjugate
The complex elements of the $N\times N$ matrix $A=\{a_{jk}:j,k=0,\dots N-1\}$ can be written as
$a_{jk}=a_{jk0}+ia_{jk1}.$ (3)
We again involve two mentioned above $n$-qubit subsystems $R$ nd $C$ and add the third one-qubit subsystem $M$ whose state is $|0\rangle_{M}$ for the real part of $a_{jk}$ and $|1\rangle_{M}$ for the imaginary part. Then the elements $a_{jk}$ of the matrix $A$ can be encoded into the superposition state of three subsystems $R$, $C$ and $M$ as follows:
$|\Psi_{A}\rangle=\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}\sum_{m=0}^{1}a_{jkm}|j\rangle_{R}|k\rangle_{C}|m\rangle_{M}.$ (4)
With such pure state, the Hermitian conjugation of $A$ can be realized by the following operator:
$W=SWAP_{RC}\sigma^{(z)}_{M},$ (5)
where $SWAP_{RC}$ exchanges states of the subsystems $R$ and $C$ (in other words, it performs transposition of $A$), and $\sigma^{(z)}_{M}$ is the $\sigma^{(z)}$-operator which is applied to $M$ and thus performs complex conjugation of $A$. Thus the matrix encoded into the state $W|\Psi_{A}\rangle$ is the Hermitian conjugate of $A$,
$|\tilde{\Psi}_{A}\rangle=W|\Psi_{A}\rangle=\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}\sum_{m=0}^{1}a_{jkm}(-1)^{m}|j\rangle_{R}|k\rangle_{C}|m\rangle_{M}.$ (6)
The Hermitian conjugation algorithm can be used as a subroutine in other algorithms, as shown in Sec.III.1. The circuit for Hermitian conjugation is presented in Fig.1.
Figure 1: The circuit for Hermitian conjugation, $Z\equiv\sigma^{(z)}$.
We shall notice that using the new encoding of the complex matrix in the superposition state of the triple system $R\cup C\cup M$ requires appropriate modifications for the algorithms in Refs.ZQKW_2024 ; ZBQKW_arXive2024 . In addition, since algorithms of determinant calculation and matrix inversion ZBQKW_arXive2024 use the row-wise matrix encoding, each row requires its own additional one-qubit subsystem to label the real and imaginary parts of its matrix elements, i.e., $N$ one-qubit subsystems $M_{i}$, $i=0,\dots,N-1$, are required. In Sec.III, we consider modification of multiplication algorithm.
II.2 Weakened normalization constraint for matrix encoding
Equality (2) represents a constraint imposed on the absolute values of matrix elements. However, we can weaken this constraint reducing it to the inequality
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}<1.$ (7)
To satisfy inequality (7), the matrix $A$ has to be properly prepared. Namely, instead of the matrix $A$ we encode the matrix $\tilde{A}=\{\tilde{a}_{ij}:i,j=0,\dots,N-1\}$,
$\tilde{A}=\frac{1}{\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}+c}A,\;\;c>0,$ (8)
where the positive parameter $c$ is conventional, it is required to guarantee that condition (7) for the matrix $\tilde{A}$ is a strict inequality. To encode the prepared matrix $\tilde{A}$, we introduce, along with the subsystems $R$ and $C$ enumerating rows and columns of $\tilde{A}$, the one-qubit subsystem $K$ and replace the initial state (1) with the following one:
$|\Psi_{A}\rangle$ $=$ $b|0\rangle_{R}|0\rangle_{C}|0\rangle_{K}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}\tilde{a}_{jk}|j\rangle_{R}|k\rangle_{C}|1\rangle_{K}$ (9)
with the normalization
$|b|^{2}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|\tilde{a}_{jk}|^{2}=1,$ (10)
where $b$ serves to satisfy the constraint (10) and can be determined as
$|b|^{2}=1-\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|\tilde{a}_{jk}|^{2}.$ (11)
II.3 Circuit for new encoding of initial state
Now we combine the modifications of the state encoding presented in Eqs.(4) and (9) and encode the elements $\tilde{a}_{ij}$ decomposed into the real and imaginary parts (3) as follows:
$|\Psi_{A}\rangle$ $=$ $(b^{(i)}_{0}|0\rangle_{M}+b^{(i)}_{1}|1\rangle_{M})|0\rangle_{R}|0\rangle_{C}|0\rangle_{K}+$ (13)
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}(\tilde{a}_{jk0}|0\rangle_{M}+\tilde{a}_{jk1}|1\rangle_{M})|j\rangle_{R}|k\rangle_{C}|1\rangle_{K},$
$|b|^{2}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|\tilde{a}_{jk}|^{2}=1,$
Replacing matrix representation (1) with representation (13) is shown in Fig.2.
Figure 2: Replacement of matrix encoding (1) with encoding (13) for organizing effective manipulations with complex matrices and weakening the constraints on the matrix elements.
Below we omit tilde over the elements of the encoded matrix for the brevity, i.e., write $a_{jk}$ instead of $\tilde{a}_{jk}$ assuming that elements $a_{jk}$ satisfy inequality (7).
III Matrix multiplication
We modify algorithm for matrix multiplication proposed in ZQKW_2024 and improved in ZBQKW_arXive2024 . Modification is based on passing from encoding (1) (used in ZQKW_2024 ) to encoding (13) for the square $N\times N$ matrices $A^{(1)}$ and $A^{(2)}$ ($A^{(i)}=\{a^{(i)}_{jk}:j,k=0,\dots,N-1\}$, $i=1,2$) according to Sec.II. We assume that the both matrices $A^{(1)}$ and $A^{(2)}$ are prepared in accordance with formula (8) and therefore their elements satisfy inequality (7).
We first introduce four $n$-qubit registers $R^{(i)}$ and $C^{(i)}$ , $i=1,2$, enumerating, respectively, the rows and columns of the matrices $A^{(i)}$, $1,2$. Then, we introduce four one-qubit subsystems $M^{(i)}$ and $K^{(i)}$, $i=1,2$. The subsystems $M^{(i)}$, $i=1,2$, are used to label the real and imaginary parts of the matrix elements of the $i$th matrix ($a^{(i)}_{jk}=a^{(i)}_{jk0}+ia^{(i)}_{jk1}$), and also to label the real and imaginary parts of the parameters $b^{(i)}$ ($b^{(i)}=b^{(i)}_{0}+ib^{(i)}_{1}$). The subsystems $K^{(i)}$, $i=1,2$, are needed to weaken the constraints imposed on the elements $a^{(i)}_{jk}$. Thus, the pure states encoding the elements of matrices $A^{(i)}$, $i=1,2$, are in form (13) with normalization (13):
$|\Psi^{(i)}\rangle$ $=$ $(b^{(i)}_{0}|0\rangle_{M^{(i)}}+b^{(i)}_{1}|1\rangle_{M^{(i)}})|0\rangle_{R^{(i)}}|0\rangle_{C^{(i)}}|0\rangle_{K^{(i)}}+$ (15)
$\sum_{j,k=0}^{N-1}(a^{(i)}_{jk0}|0\rangle_{M^{(i)}}+a^{(i)}_{jk1}|1\rangle_{M^{(i)}})|j\rangle_{R^{(i)}}|k\rangle_{C^{(i)}}|1\rangle_{K^{(i)}},$
$|b^{(i)}|^{2}+\sum_{j,k}|a^{(i)}_{jk}|^{2}=1,$
where the constant parameters $b^{(i)}$ are introduced in accordance with Sec.II.2. The initial state of the whole system reads
$|\Phi_{0}\rangle=|\Psi_{2}\rangle\otimes|\Psi_{1}\rangle=$ (16)
$\sum_{m_{1},m_{2}=0}^{1}b^{(1)}_{m_{1}}b^{(2)}_{m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}+$
$\sum_{{j_{1},k_{1},}\atop{j_{2},k_{2}=0}}^{N-1}\sum_{m_{1},m_{2}=0}^{1}\left(a^{(1)}_{j_{1}k_{1}m_{1}}a^{(2)}_{j_{2}k_{2}m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}\right)$
×$\times|j_{1}\rangle_{R_{1}}|k_{1}\rangle_{C^{(1)}}|j_{2}\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{0}\rangle,$
where we select the terms with the states $|j\rangle_{K^{(1)}}|j\rangle_{K^{(2)}}$, $j=0,1$, which are needed below, while the terms with the states $|j\rangle_{K^{(1)}}|k\rangle_{K^{(2)}}$, $j\neq k$, are collected in the garbage $|g_{0}\rangle$.
From the superposition state $|\Phi_{0}\rangle$, we have to select the terms with the states $|k\rangle_{C^{(1)}}|k\rangle_{R^{(2)}}$, $k=0,\dots,N-1$, i.e., $k_{1}=j_{2}$ in the second part in the rhs of formula (16). For that purpose we apply the CNOTs $W_{j}$ to the $j$th qubits of $C^{(1)}$ and $R^{(2)}$,
$W_{j}$ $=$ $|1\rangle_{C^{(1)}_{j}}\;{}_{C^{(1)}_{j}}\langle 1|\otimes\sigma^{(x)}_{R^{(2)}_{j}}+|0\rangle_{C^{(1)}_{j}}\;{}_{C^{(1)}_{j}}\langle 0|\otimes I_{R^{(2)}_{j}},\;\;j=1,\dots,n.$ (17)
where $\sigma^{(x)}_{R^{(2)}_{j}}$ and $I_{R^{(2)}_{j}}$ are, respectively, the $\sigma^{(x)}$-operator and identity operator applied to the $j$th qubit of the subsystem $R^{(2)}$. We have chosen the qubits of $C^{(1)}$ as control ones. All operators $W_{j}$, $j=1,\dots,n$, commute with each other. Applying the operator $W^{(0)}_{C^{(1)}R^{(2)}}$,
$W^{(0)}_{C^{(1)}R^{(2)}}=\prod_{j=1}^{n}W_{j},$ (18)
to $|\Phi_{0}\rangle$ we obtain:
$|\Phi_{1}\rangle=W^{(0)}_{C^{(1)}R^{(2)}}|\Phi_{0}\rangle=$ (19)
$\sum_{m_{1},m_{2}=0}^{1}b^{(1)}_{m_{1}}b^{(2)}_{m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}+$
$\sum_{{j_{1},j,}\atop{k_{2}=0}}^{N-1}\sum_{m_{1},m_{2}=0}^{1}\left(a^{(1)}_{j_{1}jm_{1}}a^{(2)}_{jk_{2}m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}\right)$
×$\times|j_{1}\rangle_{R^{(1)}}|j\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{1}\rangle.$
The second part in the rhs of (19) collets the terms with the state $|0\rangle_{R_{2}}$ (the other terms are garbage by definition; their are collected in $|g_{1}\rangle$ above). We note that the first part in the rhs of (19) remains unchanged after applying the operator $W^{(0)}_{C^{(1)}R^{(2)}}$.
Next, we apply the Hadamard transformations
$W^{(1)}_{C^{(1)}}=H^{\otimes n}$ (20)
to the qubits of $C^{(1)}$ and select the terms with $|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}$ putting others to the garbage $|g_{2}\rangle$:
$|\Phi_{2}\rangle=W^{(1)}_{C^{(1)}}|\Phi_{1}\rangle=$ (21)
$\frac{1}{2^{n/2}}\sum_{m_{1},m_{2}=0}^{1}b^{(1)}_{m_{1}}b^{(2)}_{m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}$
$+\frac{1}{2^{n/2}}\sum_{j_{1},k_{2}=0}^{N-1}\sum_{j=0}^{N-1}\sum_{m_{1},m_{2}=0}^{1}\left(a^{(1)}_{j_{1}jm_{1}}a^{(2)}_{jk_{2}m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}\right)$
×$\times|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{2}\rangle.$
We note that the operator $W^{(1)}_{C^{(1)}}$ transforms $|g_{1}\rangle$ as well. However, we don’t need details of that transformation. The similar statement can be referred to the operators $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ and $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}$ below.
Now, to complete multiplication, we have to organize proper combinations of products of real and imaginary parts of matrices $A^{(i)}$, $i=1,2$, to form the real and imaginary parts of the result. To this end we introduce the following control operator:
$W_{M^{(1)}M^{(2)}}=|1\rangle_{M^{(2)}}\,{}_{M^{(2)}}\langle 1|\otimes\sigma^{(x)}_{M^{(1)}}\sigma^{(z)}_{M^{(1)}}+|0\rangle_{M^{(2)}}\,{}_{M^{(2)}}\langle 0|\otimes I_{M^{(1)}}.$ (22)
We also introduce the controlled operator
$W_{K^{(1)}K^{(2)}}=|1\rangle_{K^{(1)}}\,{}_{K^{(1)}}\langle 1|\otimes\sigma^{(x)}_{K^{(2)}}+|0\rangle_{K^{(1)}}\,{}_{K^{(1)}}\langle 0|\otimes I_{K^{(2)}}$ (23)
which labels all needed terms by the state $|0\rangle_{K^{(2)}}$. We include the operators $W_{M^{(1)}M^{(2)}}$ and $W_{K^{(1)}K^{(2)}}$ into the operator
$W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}=W_{K^{(1)}K^{(2)}}H_{M^{(2)}}W_{M^{(1)}M^{(2)}},$ (24)
where $H_{M^{(2)}}$ is the set of the Hadamard operators applied to single qubit of $M^{(2)}$, and apply $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ to $|\Phi_{2}\rangle$, thus obtaining
$|\Phi_{3}\rangle=W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}|\Phi_{2}\rangle=$ (25)
$\frac{1}{2^{(n+1)/2}}\Big{(}(b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1})|0\rangle_{M^{(1)}}+(b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0})|1\rangle_{M^{(1)}}\Big{)}$
×$\times|0\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}$
$+\frac{1}{2^{(n+1)/2}}\sum_{j_{1},k_{2}=0}^{N-1}\sum_{j=0}^{N-1}\left((a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}0}-a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}1})|0\rangle_{M^{(1)}}\right.$
$\left.+(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}1}+a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}0})|1\rangle_{M^{(1)}}\right)$
×$\times|0\rangle_{M^{(2)}}|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}+|g_{3}\rangle.$
Here the second term contains the desired matrix product and all extra terms are collected in the garbage $|g_{3}\rangle$. We emphasize that the operator $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ is a novelty in comparison with multiplication algorithm in ZBQKW_arXive2024 .
Next, to label and remove the garbage, we prepare two one-qubit ancillae $B$ and $\tilde{B}$ in the ground state $|0\rangle_{B}|0\rangle_{\tilde{B}}$, introduce the projector
$P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}=|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{M^{(2)}}|0\rangle_{K^{(2)}}\;{}_{C^{(1)}}\langle 0|_{R^{(2)}}\langle 0|_{M^{(2)}}\langle 0|_{K^{(2)}}\langle 0|$ (26)
and the controlled operator $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}$,
$W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}=P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}\otimes\sigma^{(x)}_{B}\sigma^{(x)}_{\tilde{B}}$ (27)
$+(I_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}-P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}})\otimes I_{B\tilde{B}},$
of the depth $O(n)$ with $2(n+1)$-qubit control register. Applying this operator to $|\Phi_{3}\rangle|0\rangle_{B}|0\rangle_{\tilde{B}}$ we obtain
$|\Phi_{4}\rangle=W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}|\Phi_{3}\rangle|0\rangle_{B}|0\rangle_{\tilde{B}}=$ (28)
$\frac{1}{2^{(n+1)/2}}\Big{(}(b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1})|0\rangle_{M^{(1)}}+(b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0})|1\rangle_{M^{(1)}}\Big{)}$
×$\times|0\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}|1\rangle_{B}|1\rangle_{\tilde{B}}+$
$\frac{1}{2^{(n+1)/2}}\sum_{{j_{1},k_{2}}\atop{j=0}}^{N-1}\Big{(}(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}0}-a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}1})|0\rangle_{M^{(1)}}+(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}1}+a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}0})|1\rangle_{M^{(1)}}\Big{)}$
×$\times|0\rangle_{M^{(2)}}|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}\ |1\rangle_{B}|1\rangle_{\tilde{B}}+|g_{3}\rangle|0\rangle_{B}|0\rangle_{\tilde{B}}.$
Now we perform the conditional measurement FZQW_arxive2025
$W^{(4)}_{B\tilde{B}}=|1\rangle_{B}\,{}_{\tilde{B}}\langle 1|\otimes M_{\tilde{B}}+|0\rangle_{B}\,{}_{\tilde{B}}\langle 0|\otimes I_{\tilde{B}},$ (29)
where $M_{\tilde{B}}$ is the measurement operator applied to $\tilde{B}$. Then,
$|\Phi_{5}\rangle=W^{(4)}_{B\tilde{B}}|\Phi_{4}\rangle=|\Psi_{out}\rangle\,|0\rangle_{M_{2}}|0\rangle_{C_{1}}|0\rangle_{R_{2}}|0\rangle_{K_{2}}|1\rangle_{B},$ (30)
$|\Psi_{out}\rangle=G^{{-1}}\Big{(}\Big{(}(b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1})|0\rangle_{M^{(1)}}+(b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0})|1\rangle_{M^{(1)}}\Big{)}$
×$\times|0\rangle_{R^{(1)}}|0\rangle_{C^{(2)}}|0\rangle_{K_{1}}+$
$\sum_{{j_{1},k_{2},}\atop{j=0}}^{N-1}\left((a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}0}-a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}1})|0\rangle_{M^{(1)}}+(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}1}+a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}0})|1\rangle_{M^{(1)}}\right)$
×$\times|j_{1}\rangle_{R^{(1)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}\Big{)}=$
$G^{-1}\left((\hat{b}_{0}|0\rangle_{M^{(1)}}+\hat{b}_{1}|1\rangle_{M^{(1)}})|0\rangle_{R^{(1)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}\right.$
$\left.+\sum_{j,k=0}^{N-1}(\hat{a}_{jk0}|0\rangle_{M^{(1)}}+\hat{a}_{jk1}|1\rangle_{M^{(1)}})|j\rangle_{R^{(1)}}|k\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}\right),$
where $G=\left(|b^{(1)}b^{(2)}|^{2}+\sum_{j,k=0}^{N-1}\Big{|}\sum_{l=0}^{N-1}a^{(1)}_{jl}a^{(2)}_{lk}\Big{|}^{2}\right)^{1/2}$ is the normalization factor,
$\hat{b}_{0}=b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1},$ (31)
$\hat{b}_{1}=b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0},$
$\hat{a}_{jk0}=\sum_{l=0}^{N-1}(a^{(1)}_{jl0}a^{(2)}_{lk0}-a^{(1)}_{jl1}a^{(2)}_{lk1}),$
$\hat{a}_{jk1}=\sum_{l=0}^{N-1}(a^{(1)}_{jl0}a^{(2)}_{lk1}+a^{(1)}_{jl1}a^{(2)}_{lk0}),$
so that $\hat{A}=A^{(1)}A^{(2)}=\{\hat{a}_{jk}:j,k=0,\dots,N-1\}$, $\hat{a}_{jk}=\hat{a}_{jk0}+i\hat{a}_{jk1}$, $\hat{b}=\hat{b}_{0}+i\hat{b}_{1}$, where $\hat{a}_{jkm}$, $\hat{b}_{m}$, $m=0,1$, are real numbers. We note that the first term in the final expression for $|\Phi_{5}\rangle$ plays the same role as the first term in the expression for $|\Phi_{0}\rangle$ in (16). It weakens the constraint on the normalization of the elements of the resulting matrix $\hat{A}$.
Figure 3: (a) The circuit for the matrix multiplication algorithm; $|\Phi_{out}\rangle$ is formed by subsystems $R_{1}$, $C_{2}$, $K_{1}$ and $M_{1}$; $X\equiv\sigma^{(x)}$, $Z\equiv\sigma^{(z)}$; we omit subscripts in notations $W^{(j)}$ for the brevity; operator $W^{(2)}$ is a novelty in comparison with earlier algorithm in ZBQKW_arXive2024 ; (b) Notation for multi-qubit CNOT.
III.1 Manipulations with input matrices
In this section we consider several manipulations with the input matrices $A^{(1)}$ and $A^{(2)}$. Instead of multiplication $A^{(1)}A^{(2)}$ we can perform multiplications $(A^{(1)})^{\dagger}A^{(2)}$, $A^{(1)}(A^{(2)})^{\dagger}$, $(A^{(1)})^{\dagger}(A^{(2)})^{\dagger}$, $(A^{(1)}A^{(2)})^{\dagger}=(A^{(2)})^{\dagger}(A^{(1)})^{\dagger}$ and change order of the multiplier. With this purpose we introduce the following operators.
To organize the Hermitian conjugate of $A^{(i)}$, $i=1,2$, we use operator (5) introduced in Sec.II.1:
$Q^{(i)}_{R^{(i)}C^{(i)}M^{(i)}}=SWAP_{R^{(i)}C^{(i)}}\sigma^{(z)}_{M^{(i)}},\;\;i=1,2.$ (32)
To change position of matrices we apply the operator
$Q^{(3)}_{R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}}=SWAP_{R^{(1)}C^{(1)}}SWAP_{R^{(2)}C^{(2)}}SWAP_{M^{(1)}M^{(2)}}.$ (33)
The circuit implementing these operators is shown in Fig.4. For instance, to calculate the product $A^{(1)}(A^{(2)})^{\dagger}$ we apply the operator $Q^{(2)}_{R^{(2)}C^{(2)}M^{(2)}}$ to the state $|\Phi_{0}\rangle$. To calculate $(A^{(1)}A^{(2)})^{\dagger})$ we apply the operators $Q^{(3)}_{R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}}$, $Q^{(2)}_{R^{(2)}C^{(2)}M^{(2)}}$, $Q^{(1)}_{R^{(1)}C^{(1)}M^{(1)}}$ to the state $|\Phi_{0}\rangle$. Application of operators $Q^{(i)}$, $i=1,2,3$, can be controlled by additional qubits $q^{(i)}$, i.e., by the following controlled operators:
$\hat{Q}^{(i)}_{S^{(i)}}=|1\rangle_{q^{(i)}}\,{}_{q^{(i)}}\langle 1|\otimes Q^{(i)}_{S^{(i)}}+|0\rangle_{q^{(i)}}\,{}_{q^{(i)}}\langle 0|\otimes I_{S^{(i)}},$ (34)
where $S^{(i)}=R^{(i)}C^{(i)}M^{(i)}$, $i=1,2$, $S^{(3)}=R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}$ and $I_{S^{(i)}}$ is the identity operator applied to $S^{(i)}$. The state $|\varphi\rangle_{q}=|n_{1}\rangle_{q^{(1)}}|n_{2}\rangle_{q^{(2)}}|n_{3}\rangle_{q^{(3)}}$, $n_{i}=0,1$ ($i=1,2,3$) of the controlling qubits is not changed after running the algorithm.
Of cause calculating $(A^{(1)}A^{(2)})^{\dagger}$ can be dome by performing Hermitian conjugation after multiplication $A^{(1)}A^{(2)}$, this is another option.
Figure 4: Controlled manipulations with input matrices $A^{(1)}$ and $A^{(2)}$.
III.2 Measuring normalization $G$ and discussion
Although the controlled measurement allows to resolve the problem of exponential decrease of the success probability with the number of qubits encoding considered matrices we lose the information about normalization $G$. However, this information might required in obtaining the final result. Below we explore the method of indirect measuring the normalization $G$.
Let $S_{1}=|b^{(1)}b^{(2)}|^{2}$, this value is known due to the matrix preparation, see eq.(11). If, along with $S_{1}$, we know $S_{1}/G^{2}=\tilde{S}_{1}$, then $G=\sqrt{S_{1}/\tilde{S}_{1}}$. To find $\tilde{S}_{1}$, we refer to the state $|\Phi_{5}\rangle$ (30) and measure the state of the subsystem $K_{1}$ with the output 0. It is easy to understand that the probability of this output equals $\tilde{S}_{1}$. and can be found in result of multiple runs of the algorithm.
It is important that both introduced extensions do not decline the parameters of the multiplication circuit such as depth and space. Fig.3 shows that the depth is determined by the operator $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}$, while manipulations in Sec.III.1, see Fig.4, increase this depth only by insignificant constant factor. Thus, the estimation of the depth remains $O(\log N)$ similar to the estimation in ZBQKW_arXive2024 . Regarding the space, the circuit in Fig.3 includes only five additional one-qubit subsystems ($K^{(i)}$, $M^{(i)}$, $i=1,2$ , and $\tilde{B}$) in comparison with the circuit in ZBQKW_arXive2024 with one of them (the ancilla $\tilde{B}$) is needed for organization of controlled measurement. Therefore, the space estimation remains $O(\log N)$ as well.
We have to note that the above measurement destroys the part of superposition quantum state in which the result of matrix multiplication is stored. Therefore, to get the full information, i.e., normalization $G$ and matrix product, we have to use the set of runs to probabilistically determine $G$ and then run the algorithm one more time to end up with the state $|\Phi_{6}\rangle$ encoding the product $A^{(1)}A^{(2)}$. However, if we don’t need $G$, then the single run is enough.
IV Conclusion
In this paper we continue studying matrix-manipulation algorithms that are proposed in ZQKW_2024 ; ZBQKW_arXive2024 and based on the encoding the matrix elements into the pure superposition state of a quantum system. Although those algorithms work with complex numbers in general, the complex conjugation (in particular, Hermitian conjugation) can not be realized in frames of those settings. This disadvantage is removed by the first our extension which consists in separating the real and imaginary parts of the complex matrix under consideration by encoding these two parts into different probability amplitudes of a pure superposition state. Such extension allows constructing effective subroutine for operations typical for complex numbers, which are complex and Hermitian conjugations. However, this separation of real and imaginary parts requires proper modifications of the quantum algorithms for matrix addition, multiplication, determinant calculation, matrix inversion and solving linear systems that are described in ZQKW_2024 ; ZBQKW_arXive2024 . In our paper, we present such modification for the multiplication algorithm and demonstrate manipulation with the input matrices using Hermitian conjugation subroutine.
Another disadvantage of the algorithms in Refs.ZQKW_2024 ; ZBQKW_arXive2024 is normalization constraint which is generated by the normalization condition for the superposition pure quantum state. This requirement is weakened by the second our extension which adds an additional term in the superposition state encoding a particular matrix. This additional term leads to replacing the normalization equality (2) with inequality (7) that imposes only the upper boundary on the sum of the squared absolute values of the matrix elements.
Of course, each of these modifications requires involving an extra qubit. Therefore, along with two subsystems $R^{(i)}$ and $C^{(i)}$, $i=1,2$, enumerating rows and columns of the matrices $A^{(i)}$, $i=1,2$, the $i$th matrix is associated with the one-qubit subsystem $M^{(i)}$, responsible for separate encoding of the real and imaginary parts of complex matrix elements, and the one-qubit subsystem $K^{(i)}$ allowing to introduce an extra term into the superposition state encoding the matrix $A^{(i)}$, namely the presence of this extra term weakens the normalization constraint. Thus, only four additional qubits appear in the modified multiplication algorithm. We note that the second one-qubit ancilae $\tilde{B}$ is needed for organization the controlled measurement which removes the problem of small success probability that appears in the case of direct uncontrolled measurement of the ancilla stat.
We emphasize that both extensions do not significantly decline the depth and space of the multiplication algorithm. Estimations for both of them remain $O(\log N)$, which is similar to the appropriate characteristics of the multiplication algorithm in ZBQKW_arXive2024 .
Similar modifications can be included in other matrix-manipulation algorithms considered in Refs.ZQKW_2024 ; ZBQKW_arXive2024 which is postponed for future study.
Acknowledgments. This project is supported by the National Natural Science Foundation of China (Grants No. 12031004, No. 12271474 and No. 61877054). The work was also partially funded by a state task of Russian Fundamental Investigations (State Registration No. 124013000760-0).