### v4

# Controlled measurement, Hermitian conjugation and normalization in matrix-manipulation algorithms

Abstract
In this paper, we solve three important problems that are revealed, in particular, in matrix-manipulation algorithms. The principal novelty is introducing the concept of controlled measurement that solves the problem of small access probability to the desired state of ancilla and possesses several remarkable properties. We also introduce separate encoding of the real and imaginary parts of a complex matrix that allows to include the Hermitian conjugation into the list of matrix manipulations. Finally, we weaken the constraints on the modulus of matrix elements unavoidably imposed by the normalization condition for a pure quantum state. The controlled measurement together with both other extensions are implemented into the matrix multiplication algorithm. The appropriate circuits are presented.
I Introduction
Quantum algorithms represent a rapidly developing area of quantum informatics whose progress is stimulated by prospective enormous advantages of quantum algorithms over their classical counterparts. We have to acknowledge such famous algorithms as the Deutsch algorithm Deu demonstrating effect of quantum parallelism in operation with Boolean functions, Shor’s factoring algorithm Sho1 ; Sho2 utilising the quantum Fourier transform QFT1 ; QFT2 ; QFT3 , Grover’s search algorithm Gro , phase-estimation algorithm for the Hermitian QFT3 and non-Hermitian Wang matrices .
We consider a particular application area of quantum algorithms which covers matrix algebraic operations, such as multiplication, addition, determinant calculation, matrix inversion and its application to linear system solver. During recent years, there are many quantum algorithms solving various aspects of linear algebra. Among them are HHL-algorithm for solving systems of linear equations HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 which was applied, in particular, to Gaussian process regression ZhaoZ , algorithm for solving linear differential equations Berry , algorithm for evaluation of matrix functions Tak .
A possible approach to quantum algorithms for matrix manipulations is proposed in ZZRF . This approach refers to matrices as operators to be applied to some vector. Therefore, it is essential in that approach that arbitrary nonunitary and non-Hermitian matrix must be transferred to the unitary one before been included into the quantum algorithm. Thus, first of all, one has to replace a given matrix $A$ (rectangular in general) with the Hermitian one and then transfer it to the unitary operator by exponentiating which can be approximated via Trotterization Trotter ; Suzuki ; T ; AT and Baker-Champbell-Hausdorff method Feldman ; BC . To implement such matrix transformations effectively, the set of special matrices is proposed to be prepared in advance. The phase estimation subroutine QFT3 is included into that algorithm.
However, mentioned algorithms do not cover all achievement in matrix-manipulation techniques. A special block-encoding model GSLW is another approach for implementing matrix algebra. Different type of block-encoding was used in the algorithm in Ref.DFZ_2020 to embed the inverse of the matrix into the unitary transformation. In LWWZ , matrix multiplication was performed trough the binary encoding of matrix elements into the pure states of quantum systems with subsequent binary multiplication via Toffoli gates. Matrix multiplication over some rings, in particular, multiplication of Boolean matrices, was studied in KN .
Some algorithms for matrix inversion have been also developed. As the first one we call the mentioned above HHL algorithm HHL where the inversion was used for solving a linear system. However, the problem of inversion of matrix eigenvalues has not been resolved in frames of quantum algorithm therein. The matrix inversion algorithm based on the singular value decomposition is presented in MRTC ; NJ . In both papers the authors apply the function evaluation problem GSLW to approximate the inverse of singular values by odd-polynomial. This approach requires introducing a special scale parameter depending on the interval where all nonzero singular values are distributed.
Recently the alternative matrix-manipulation algorithms based on the encoding the matrix elements into the complex probability amplitudes of a pure superposition state of a certain subsystem have been developed QZKW_arxive2022 ; ZQKW_2024 ; ZBQKW_arXive2024 . The principal difference of such approach is that a matrix is considered as a part of the quantum state of a system rather than a unitary operator applied to this system. Therefore the elements of the resulting matrix (obtained after addition or multiplication of other matrices, or after matrix inversion) can be measured up to some constant factor. The algorithms for matrix manipulations, such as inner product, sum and product of two matrices, calculation of determinant, matrix inversion, and solving linear systems, based on special unitary operators applied to the quantum system was proposed in QZKW_arxive2022 . The realization of those unitary operators in terms of the basic quantum operations was explored later in ZQKW_2024 (inner product, matrix addition and multiplication) and then in ZBQKW_arXive2024 (calculating determinant and inverse matrix, linear system solver). Those unitary operators are based on the multi-qubit controlled operators and are supplemented by the one-qubit ancilla measurement at the final stage of the algorithm for selecting the useful result and removing the garbage. The major disadvantage of those algorithms is the small probability of obtaining the required ancilla state after measurement (small success probability). This probability decreases exponentially with the dimension of the considered matrix thus reducing efficiency of the matrix-manipulation algorithms.
The principal novelty of our paper is that we replace the above measurement of the one-qubit ancilla state with the so-called controlled measurement, i.e., the measurement controlled by the quantum state of another one-qubit ancilla. In other words, we replace the one-qubit ancilla with the two-qubit one so that the state of the first qubit governs the measurement operator applied to the second qubit. This formally simple but crucial modification removes the above disadvantage of small success probability. Unlike the algorithm implementing the usual measurement, which requires multiple running to reach the desired ancilla state, the algorithm involving the controlled measurement assumes the single run to achieve the goal. This is the main advantage of the new algorithm. Such controlled measurement can be included into all the algorithms in Refs.ZQKW_2024 ; ZBQKW_arXive2024 as a subroutine replacing the usual ancilla measurement and thus significantly increasing their relevance. We note that the controlled measurement has been implemented in Ref. ZBQKW_arXive2024 without full description. In our paper, we present the detailed analysis of this concept. The phenomenon of quantum control of classical operations (measurement), i.e., the control by the quantum state of the ancilla qubit, is a consequence of the deep relation between the quantum and classical physics and must exist relying on reality of quantum superposition states. The demonstrated advantage of the controlled measurement over the usual measurement motivates research on its particular realization in terms of acknowledged quantum and classical operators. The applicability of controlled measurement can be extended to all measurement based quantum algorithms where the success probability is of principal meaning. This refers, to the measurement-based quantum algorithms WWLN ; ZDB ; ZBD ; Wei ; WeiBook , in particular, to the HHL-algorithm for solving linear systems HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 , various applications of measurement in quantum algorithms are reviewed in USBGSN .
Along with implementing the controlled measurement, we also introduce two following extensions to the matrix-manipulation algorithms. (i) Although those algorithms deal with complex matrices in general, some operations (including Hermitian conjugation) can not be realized in frames of those settings and therefore certain constraints are to be imposed on the variety of matrix manipulations. To remove this disadvantage and construct more flexible algorithms for matrix algebra we propose to separate the real and imaginary parts of a matrix by encoding them into two orthogonal subspaces of the whole state space. This separation can be done by involving additional one-qubit subsystem whose state serves to label the real and imaginary parts of matrix elements. As a consequence, all probability amplitudes are real in such representation. We notice that the Hermitian conjugation is not a unitary operator (similar to other matrix-manipulations in ZQKW_2024 ; ZBQKW_arXive2024 ), whose realization becomes possible due to its application to the matrix encoded into the quantum state in a specific way. (ii) Another issue addressed in this paper is weakening the normalization constraint for the elements of the matrices encoded into quantum states. We recall that such constraint is imposed by the normalization condition for a pure quantum state. For that purpose we introduce an additional probability amplitude into the superposition state encoding considered matrix. This can be done involving one more auxiliary one-qubit subsystem.
Both of the above modifications are applicable to the algorithms involving encoding of particular quantum states. For instance, the already mentioned HHL-algorithm HHL uses encoding the known right-hand side of the linear system, the matrix-manipulation algorithms in ZZRF encode the vectors subjected to the matrix operations, the algorithms in ZQKW_2024 ; ZBQKW_arXive2024 encode the matrices for calculating addition, multiplication, inversion, quantum machine learning BWPRWL ; A ; RML requires encoding the input data, encoding the large data sets is utilized in the least-square linear-regression algorithms WBL ; SSP ; Wang2 and so on.
Of course, the controlled measurement and both above extensions require modifications of the mentioned algorithms, in particular, the algorithms proposed in ZQKW_2024 ; ZBQKW_arXive2024 . Below we present the detailed analysis of the appropriate modifications introduced into the multiplication algorithm (see appendix in ZBQKW_arXive2024 ) and demonstrate manipulations with the input matrices involving Hermitian conjugation. However, similar modifications can be implemented in other matrix-manipulation algorithms such as inner product, addition, determinant calculation, matrix inversion, linear system solving. All quantum circuits associated with controlled measurement and two above extensions, as well as the circuit for modified multiplication algorithm, are included.
The paper is organized as follows. The controlled measurement is proposed in Sec.II with all necessary details. In Sec.III, we consider applications of controlled measurements to algorithms of matrix manipulations. Then, in Sec.IV, we introduced two pointed above extensions to the matrix-manipulation algorithms. In Sec.V, we implement the controlled measurement together with two above extensions into the matrix multiplication algorithm. In the same section, we demonstrate possibility of manipulating the input matrices using subroutine of Hermitian conjugation and show how to measure the output normalization constant that is lost because of the controlled measurement of the ancilla state at the last stage of the algorithm. Conclusions are given in Sec.VI. The particular example of matrix multiplication is considered in the Appendix, Sec.VII.
II Controlled measurement in quantum algorithms
II.1 Motivation
As mentioned in the Introduction, the necessity for introducing the controlled measurement instead of the usual one is motivated by small success probability in the ancilla-state measurement that appears, for instance, in quantum algorithms for matrix manipulations ZQKW_2024 ; ZBQKW_arXive2024 . Small success probability requires multiple runs of the algorithm to get the desired result of measurement. The controlled measurement completely removes this obstacle to effective application of the above algorithms.
As a simple example, let us consider the superposition quantum state of $n^{(S)}$ qubit subsystem $S$ and $n^{(R)}$ qubit subsystem $R$. The state of the whole system $S\otimes R$ reads
$|\Phi_{0}\rangle=\sum_{i=0}^{N^{(S)}-1}\sum_{j=0}^{N^{(R)}-1}a_{ij}|i\rangle_{S}|j\rangle_{R},\;\;\sum_{i=0}^{N^{(S)}-1}\sum_{j=0}^{N^{(R)}-1}|a_{ij}|^{2}=1,\;\;N^{(S)}=2^{n^{(S)}},\;\;N^{(R)}=2^{n^{(R)}}.$ (1)
Our purpose is to extract the (normalized) sum of $a_{ij}$ over the second subscript, i.e. we need to construct the state
$\sim\sum_{i=0}^{N^{(S)}-1}\left(\sum_{j=0}^{N^{(R)}-1}a_{ij}\right)|i\rangle_{S}.$ (2)
This task, for instance, is typical for the algorithms proposed in ZQKW_2024 ; ZBQKW_arXive2024 and therefore deserves the special consideration.
To transfer state in Eq.(1) to state in Eq.(2), we, first, apply the Hadamard transformation $H$ to each qubit of $R$ collecting all $H$ in the operator $H_{R}=H^{\otimes N^{(R)}}$:
$|\Phi_{1}\rangle=H_{R}|\Phi_{0}\rangle=\frac{1}{2^{n^{(R)}/2}}\sum_{i=0}^{N^{(S)}-1}\sum_{j=0}^{N^{(R)}-1}a_{ij}|i\rangle_{S}|0\rangle_{R}+|g\rangle_{SR}.$ (3)
Hereafter the subscript at the operator means the subsystem to which this operator is applied. All needed terms are collected in the first part of Eq.(3), while all the rest terms are collected in the garbage unnormalized state $|g\rangle_{SR}$, orthogonal to the basis states in the first term of $|\Phi_{1}\rangle$. Now, to label the garbage, we introduce the one-qubit ancilla $B_{1}$ in the ground state and apply the controlled operator
$W^{(1)}_{RB_{1}}=|0\rangle_{R}\,{}_{R}\langle 0|\otimes\sigma^{(x)}_{B_{1}}+(I_{R}-|0\rangle_{R}\,{}_{R}\langle 0|)\otimes I_{B_{1}}$ (4)
to the state $|\Phi_{1}\rangle|0\rangle_{B_{1}}$:
$|\Phi_{2}\rangle=W^{(1)}_{RB_{1}}|\Phi_{1}\rangle|0\rangle_{B_{1}}=\frac{1}{2^{n^{(R)}/2}}\sum_{i=0}^{N^{(S)}-1}\sum_{j=0}^{N^{(R)}-1}a_{ij}|i\rangle_{S}|0\rangle_{R}|1\rangle_{B_{1}}+|g\rangle_{SR}|0\rangle_{B_{1}}.$ (5)
Finally, we can remove the garbage via ancilla measurement with the desired output $|1\rangle_{B_{1}}$ thus obtaining the state
$|\Phi_{3}\rangle=|\Psi_{out}\rangle_{S}|0\rangle_{R}|0\rangle_{B_{1}},\;\;$ (6)
$|\Psi_{out}\rangle_{S}=G^{-1}\sum_{i=0}^{N^{(S)}-1}\sum_{j=0}^{N^{(R)}-1}a_{ij}|i\rangle_{S},\;\;G=\sqrt{\sum_{i=0}^{N^{(S)}-1}\left|\sum_{j=0}^{N^{(R)}-1}a_{ij}\right|^{2}}.$ (7)
However, the probability of obtaining the desired state $|1\rangle_{B_{1}}$ of the ancilla $B_{1}$ as the result of ancilla measurement (success probability) is $p=\frac{G^{2}}{2^{n^{(R)}}}\sim\frac{1}{2^{n^{(R)}}}$ and decreases exponentially with $n^{(R)}$. This means that the number of runs needed to access the desired ancilla state can be estimated as $O\left(\frac{2^{n^{(R)}}}{G^{2}}\right)$. We also notice that, if the only required result is the state $|\Phi_{3}\rangle$, then the single successful measurement of the ancilla state is enough to end up with this state of the subsystem $S\otimes R$. However, if we also need to calculate the normalization $G$, then the single successful measurement is not enough because $G$ can be calculated only probabilistically and the set of runs is requires to determine the success probability $p$, so that $G=2^{n^{(R)}/2}\sqrt{p}$.
Thus, in general, the proposed algorithm of selecting the needed state $|\Phi_{3}\rangle$ becomes ineffective requiring the large number of running to succeed in measuring the desired state $|1\rangle_{B_{1}}$. This statement motivates the search for methods allowing to overcome this obstacle removing the problem of small success probability. Such method is proposed in the next section, Sec.II.2.
II.2 Controlled measurement and its features
In this section we extend the concept of measurement as a tool allowing to get information from the quantum state of some system. First we recall the concept of measurement restricting ourselves to the one-qubit measurements NCh . The usage of measurement in our manuscript is in removing the garbage states, i.e., the states which appear as a byproduct of a quantum algorithm, see, for instance, HHL . Then we introduce the one-qubit controlled measurement as a tool allowing to avoid the small success probability revealed in Sec.II.1, thus removing the necessity of multiple running of the algorithm with the purpose to register the desired ancilla state. As a consequence, the controlled measurement makes the garbage removing much more effective. Thus, the controlled measurements becomes the preferable method for garbage removing and its application is demonstrated in Secs.III and V.
II.2.1 One-qubit measurement
Hereafter, by the measurement (projective measurement) of the one-qubit system $B_{1}$ we call the operator $M_{B_{1}}$ that, acting on the superposition of basis states
$|\Phi_{0}\rangle_{B_{1}}=\alpha|1\rangle_{B_{1}}+\beta|0\rangle_{B_{1}},\;\;|\alpha|^{2}+|\beta|^{2}=1,$ (8)
transfers it to one of the basis states, either $|1\rangle_{B_{1}}$ or $|0\rangle_{B_{2}}$ (up to the phase factor), with the probability, respectively, $|\alpha|^{2}$ and $|\beta|^{2}$, i.e.,
$M_{B_{1}}|\Phi_{0}\rangle_{B_{1}}=\left\{\begin{array}[]{ll}\displaystyle\frac{\alpha}{|\alpha|}|1\rangle_{B_{1}},&{\mbox{probability}}\;\;|\alpha|^{2}\cr\displaystyle\frac{\beta}{|\beta|}|0\rangle_{B_{1}},&{\mbox{probability}}\;\;|\beta|^{2}\end{array}\right.$. (11)
Thus, the qubit after measurement can not remain in a superposition state. The property of the measurement operator to select a particular basis states serves to remove the extra terms (called garbage) from the above quantum state, given in Eq.(5), reducing it to Eq.(6), see also HHL ; ZQKW_2024 ; ZBQKW_arXive2024 .
II.2.2 Controlled measurement
Now, along with the system $B_{1}$, we introduce another system $B_{2}$ in the ground state $|0\rangle_{B_{2}}$ and apply the operator
$W^{(2)}_{B_{1}B_{2}}=|1\rangle_{B_{1}}\,{}_{B_{1}}\langle 1|\otimes\sigma^{(x)}_{B_{2}}+|0\rangle_{B_{1}}\,{}_{B_{1}}\langle 0|\otimes I_{B_{2}}.$ (12)
to the state $|\Phi_{0}\rangle_{B_{1}}|0\rangle_{B_{2}}$:
$|\Phi_{1}\rangle_{B_{1}B_{2}}=W^{(2)}_{B_{1}B_{2}}|\Phi_{0}\rangle_{B_{1}}|0\rangle_{B_{2}}=\alpha|1\rangle_{B_{1}}|1\rangle_{B_{2}}+\beta|0\rangle_{B_{1}}|0\rangle_{B_{2}},$ (13)
thus doubling the state of $B_{1}$. Next, we introduce the following operator, called the controlled measurement, using $B_{1}$ as the controlling qubit and $B_{2}$ as the controlled qubit:
$W^{(3)}_{B_{1}B_{2}}=|1\rangle_{B_{1}}\,{}_{B_{1}}\langle 1|\otimes M_{B_{2}}+|0\rangle_{B_{1}}\,{}_{B_{1}}\langle 0|\otimes I_{B_{2}},$ (14)
where $M_{B_{2}}$ is the measurement operator, discussed in Sec.II.2.1, applied to the ancilla $B_{2}$. Notice that the operator $W^{(3)}_{B_{1}B_{2}}$ is not a unitary one due to the measurement $M_{B_{2}}$. Applying this operator to $|\Phi_{1}\rangle_{B_{1}B_{2}}$ we obtain the state
$W^{(3)}_{B_{1}B_{2}}|\Phi_{1}\rangle_{B_{1}B_{2}}=\alpha M_{B_{2}}|1\rangle_{B_{1}}|1\rangle_{B_{2}}+\beta|0\rangle_{B_{1}}|0\rangle_{B_{2}}.$ (15)
Eq.(15) demonstrates that the measurement of $B_{2}$ is switched on by the state $|1\rangle_{B_{1}}$ of $B_{1}$ and the operator $M_{B_{2}}$ is applied to the desired state $|1\rangle_{B_{2}}$. Therefore, if $\alpha\neq 0$, the measurement yields only $|1\rangle_{B_{2}}$ without the choice between $|0\rangle_{B_{2}}$ and $|1\rangle_{B_{2}}$. In other words, we deal with the perfect success in getting the desired state $|1\rangle_{B_{2}}$. This means that the second term must disappear from (15) because it includes state $|0\rangle_{B_{2}}$ rather than $|1\rangle_{B_{2}}$. Thus
$W^{(3)}_{B_{1}B_{2}}|\Phi_{1}\rangle_{B_{1}B_{2}}=\left\{\begin{array}[]{ll}\displaystyle\frac{\alpha}{|\alpha|}|1\rangle_{B_{1}}|1\rangle_{B_{2}},&\alpha\neq 0\cr\displaystyle\frac{\beta}{|\beta|}|0\rangle_{B_{1}}|0\rangle_{B_{2}},&\alpha=0\end{array}\right.$. (18)
Similarly, the controlled measurement can be applied to the state
$|\tilde{\Phi}_{1}\rangle=\alpha|E_{1}\rangle_{S}|1\rangle_{B_{1}}|1\rangle_{B_{2}}+\beta|E_{2}\rangle_{S}|0\rangle_{B_{1}}|0\rangle_{B_{2}},\;\;|\alpha|^{2}+|\beta|^{2},$ (19)
where $|E_{i}\rangle_{S}$, $i=1,2$, are some normalized states of the system $S$, yielding the following result:
$W^{(3)}_{B_{1}B_{2}}|\tilde{\Phi}_{1}\rangle=\left\{\begin{array}[]{ll}\displaystyle\frac{\alpha}{|\alpha|}|E_{1}\rangle_{S}|1\rangle_{B_{1}}|1\rangle_{B_{2}},&\alpha\neq 0\cr\displaystyle\frac{\beta}{|\beta|}|E_{2}\rangle_{S}|0\rangle_{B_{1}}|0\rangle_{B_{2}},&\alpha=0\end{array}\right.$. (22)
The controlled measurement has one interesting feature. To reveal it, we apply the controlled measurement $W^{(3)}_{B_{1}B_{2}}$ to the state $|\Phi_{1}\rangle_{B_{1}B_{2}}=|\Phi_{0}\rangle_{B_{1}}|0\rangle_{B_{2}}$. Instead of (15), we have
$W^{(3)}_{B_{1}B_{2}}|\Phi_{1}\rangle_{B_{1}B_{2}}=\alpha M_{B_{2}}|1\rangle_{B_{1}}|0\rangle_{B_{2}}+\beta|0\rangle_{B_{1}}|0\rangle_{B_{2}}.$ (23)
Again, the measurement is initiated by the first term of Eq.(23) and fixes the state $|0\rangle_{B_{2}}$ of the qubit $B_{2}$. But the same state $|0\rangle_{B_{2}}$ appears in the second term of Eq.(23). Consequently, this term also contributes to the resulting state (we emphasize that the states of $B_{2}$ in the first and second terms of the state $|\Phi_{1}\rangle_{B_{1}B_{2}}$ in Eq.(13) are orthogonal to each other, so that the second term of Eq.(13) does not contribute to the output in Eq.(18) after the controlled measurement if $\alpha\neq 0$). Therefore, the result of controlled measurement is
$W^{(3)}_{B_{1}B_{2}}|\Phi_{1}\rangle_{B_{1}B_{2}}=\Big(|1\rangle_{B_{1}}+\beta|0\rangle_{B_{1}}\Big)|0\rangle_{B_{2}}.$ (24)
This is the case when the controlled measurement is equivalent to the usual one.
Formulae (18) and (24) can be generalized as follows. We assume that the controlled state of $B_{2}$ is a superposition state, i.e., we replace the state $|\Phi_{1}\rangle_{B_{1}B_{2}}$ given in Eq.(13) with the following one:
$|\Phi_{1}\rangle_{B_{1}B_{2}}=\alpha|1\rangle_{B_{1}}(\alpha_{1}|0\rangle_{B_{2}}+\beta_{1}|1\rangle_{B_{2}})+\beta|0\rangle_{B_{1}}(\alpha_{2}|0\rangle_{B_{2}}+\beta_{2}|1\rangle_{B_{2}}),\;\;|\alpha_{i}|^{2}+|\beta_{i}|^{2}=1,\;\;i=1,2.$ (25)
We also assume that the superposition states of $B_{2}$ used in the first and second terms of Eq.(25) are not orthogonal to each other in general. Instead of Eq.(14), we have
$W^{(3)}_{B_{1}B_{2}}|\Phi_{1}\rangle_{B_{1}B_{2}}=\alpha|1\rangle_{B_{1}}M_{B_{2}}(\alpha_{1}|0\rangle_{B_{2}}+\beta_{1}|1\rangle_{B_{2}})+\beta|0\rangle_{B_{1}}(\alpha_{2}|0\rangle_{B_{2}}+\beta_{2}|1\rangle_{B_{2}}).$ (26)
Then, the measurement $M_{B_{2}}$ initiated by the first term in (26) yields the state $|0\rangle_{B_{2}}$ with the probability $|\alpha_{1}|^{2}$ and the state $|1\rangle_{B_{2}}$ with the probability $|\beta_{1}|^{2}$ (because the measurements is applied to the state of $B_{2}$ associated with the first term of Eq.(26)), both these probabilities do not depend on $\beta$. Now, to correctly write the resulting state after applying $W^{(3)}_{B_{1}B_{2}}$, we emphasize that the state $|1\rangle_{B_{1}}$ of $B_{1}$ just switches on the measurement $M_{B_{2}}$ thus destroying the superposition state of $B_{2}$ putting it to the basis state, either $|0\rangle_{B_{2}}$ or $|1\rangle_{B_{2}}$. However, the fact that the state becomes fixed holds for both terms of Eq.(26) (similar to Eq.(24)) because the qubit $B_{2}$ can not remain in a superposition state after measurement. Thus, since both terms in (26) contain two basis states $|0\rangle_{B_{2}}$ and $|1\rangle_{B_{2}}$, both of these terms yield contributions to the resulting state of quantum system after the controlled measurement. Therefore, we have
$W^{(3)}_{B_{1}B_{2}}|\Psi_{1}\rangle_{B_{1}B_{2}}=\left\{\begin{array}[]{ll}G_{1}^{-1}\Big(\alpha\alpha_{1}|1\rangle_{B_{1}}+\beta\alpha_{2}|0\rangle_{B_{1}}\Big)|0\rangle_{B_{2}},&\alpha\neq 0,\;\;{\mbox{if}}\;\;\displaystyle M_{2}|\chi_{1}\rangle_{B_{2}}=\frac{\alpha_{1}}{|\alpha_{1}|}|0\rangle_{B_{2}}\cr G_{2}^{-1}\Big(\alpha\beta_{1}|1\rangle_{B_{1}}+\beta\beta_{2}|0\rangle_{B_{1}}\Big)|1\rangle_{B_{2}},&\alpha\neq 0,\;\;{\mbox{if}}\;\;\displaystyle M_{2}|\chi_{1}\rangle_{B_{2}}=\frac{\beta_{1}}{|\beta_{1}||1\rangle_{B_{2}}}\cr e^{i\varphi_{\beta}}|0\rangle_{B_{1}}(\alpha_{2}|0\rangle_{B_{2}}+\beta_{2}|1\rangle_{B_{2}})\equiv|\Psi_{1}\rangle_{B_{1}B_{2}},&\alpha=0,\;\;\beta=e^{i\varphi_{\beta}},\;\;\varphi_{\beta}\in\mathbb{R}\end{array}\right$., (30)
where $G_{1}=\sqrt{|\alpha\alpha_{1}|^{2}+|\beta\alpha_{2}|^{2}}$, $G_{2}=\sqrt{|\alpha\beta_{1}|^{2}+|\beta\beta_{2}|^{2}}$. Thus, if $\alpha\neq 0$, we have a superposition state. We note that applying the usual measurement to the state $|\Psi_{1}\rangle_{B_{1}B_{2}}$ yields the same states as in the first and second lines in the right-hand side of Eq.(30), but the probabilities of obtaining these states are, respectively, $G_{1}^{2}=|\alpha\alpha_{1}|^{2}+|\beta\alpha_{2}|^{2}$ and $G_{2}^{2}=|\alpha\beta_{1}|^{2}+|\beta\beta_{2}|^{2}$ and equal to the probabilities of measuring the states $|0\rangle_{B_{2}}$ and $|1\rangle_{B_{2}}$ respectively. These probabilities differ from those obtained in the case of the controlled measurement. In particular, they depend on $\beta$. If $\alpha=0$, then the operator $W^{(3)}_{B_{1}B_{2}}$ does not apply the measurement $M_{B_{2}}$ as shown in the third line in the right-hand side of Eq.(30).
II.3 Properties of controlled measurement
We collect some most important properties of the controlled measurement
  1. 1.
    To access the desired state of the ancilla via the usual measurement the multiple run of the algorithm is required, whereas the single run is enough to do the same via the controlled measurement. As a consequence, the controlled measurement does not supply any classical information, unlike the ordinary measurement. In fact, we do not obtain the normalization $G$ in Eq.(7) in terms of the probability $p$, because the result of measurement of the ancilla $B_{2}$ is predictable and equals $|1\rangle_{B_{2}}$. Thus, the quantum state obtained after the controlled measurement can not be considered as a result of post-selection. We see that the controlled measurement crucially differs from the usual measurement.
  2. 2.
    The controlled measurement $W^{(3)}_{B_{1}B_{2}}$, defined in Eq.(14), is not a unitary operator due to the presence of measurement $M_{B_{2}}$. In addition, the definition of the controlled measurement, given in Eq.(14), can be generalized as follows:
    $W^{(3)}_{B_{1}B_{2}}=|1\rangle_{B_{1}}\,{}_{B_{1}}\langle 1|\otimes M_{B_{2}}+|0\rangle_{B_{1}}\,{}_{B_{1}}\langle 0|\otimes U_{B_{2}},$ (31)
    where $U_{B_{2}}$ is any unitary operator applied to $B_{2}$. This operator plays no role if the superposition state includes the state $|1\rangle_{B_{1}}$.
  3. 3.
    The controlled measurement maps the pure state into another pure state, therefore it is a trace-preserving positive map.
  4. 4.
    Although the controlled measurement is seemingly equivalent to the amplification of the probability of obtaining the desired state, it is not a variant/generalization of the amplitude amplification algorithm based on the Grover’s search algorithm NCh ; BHMP ; KB ; Liu_2024 ; KS_2025 ; CZS . The principal difference is including the measurement in the operator of controlled measurement. Therefore we deal with the operator combining both quantum and classical features. Unlike the amplitude amplification via the modifications of the Grover’s algorithm, which require multiple applying of certain unitary operator, the controlled measurement assumes the single run.
  5. 5.
    Application of the controlled measurement to the density matrix is also possible. In this case, the density matrix might be considered as a partial trace from the purified quantum state. Details are left for further study.
We also note that including the controlled measurement requires only one-qubit additional ancilla $B_{2}$ and therefore can not be considered as a space-consuming operator.
II.4 On realizability of controlled measurement
There is no principal objections to realizability of the controlled measurement. Due to the presence of the state $|1\rangle_{B_{1}}$ (which switches on the measurement $M_{B_{2}}$ of the state $|1\rangle_{B_{2}}$) in the superposition state $|\Phi_{1}\rangle_{B_{1}B_{2}}$ (see Eqs.(13), (25)), we do not need to repeat running the algorithm even if the probability of access to the desired state $|1\rangle_{B_{2}}$ via the usual measurement is small. This is the main difference between the controlled measurement and the usual measurement where the value of access probability to the desired ancilla-state has the crucial meaning. In other words, we switch on the measurement (that can be called a classical operation) by the quantum tool (via the state $|1\rangle_{B_{1}}$) relying on reality of quantum superposition states. Therefore, realizability of controlled measurement would additionally confirm the existence of such states.
Thus, the controlled measurement can be considered as applying the classical operator (measurement) to some subsystem (ancilla $B_{2}$) controlled by the quantum state of another subsystem (ancilla $B_{1}$). Notice that the opposite control (i.e., applying the quantum operator controlled by the classical state of some subsystem, i.e., by the state after measurement) is an acknowledged controlled operation used, for instance, in the teleportation algorithm BBCJPW ; NCh , this is the classical-quantum control. The quantum-quantum control is also well known, C-NOT is the simplest representative of such operation. In this case both controlling and controlled subsystems and controlled operator are quantum ones. Therefore, realization of the controlled measurement would be an example of the third type of control in the hybrid quantum-classical system and can be referred to as the quantum-classical control. At the moment, we can not suggest a particular realization of controlled measurement in terms of the well-known quantum and classical operators. However, its advantage over the usual measurement demonstrated above stimulates the search for methods realizing such operator.
III Applications of controlled measurement
III.1 Controlled measurement in algorithm developed in Sec.II.1
Let us demonstrate application of the controlled measurement in the algorithm developed in Sec.II.1. We include the one more one-qubit ancilla $B_{2}$ in the ground state and apply C-NOT $W^{(2)}_{B_{1}B_{2}}$ , given in Eq.(12), to $|\Phi_{2}\rangle|0\rangle_{B_{2}}$ (with $|\Phi_{2}\rangle$ defined in Eq.(5)) thus doubling the state of the ancilla $B_{1}$:
$|\Phi_{3}\rangle=W^{(2)}_{B_{1}B_{2}}|\Phi_{2}\rangle|0\rangle_{B_{2}}=\frac{1}{2^{n^{(R)}/2}}\sum_{i=0}^{N^{(S)}-1}\sum_{j=0}^{N^{(R)}-1}a_{ij}|i\rangle_{S}|0\rangle_{R}|1\rangle_{B_{1}}|1\rangle_{B_{2}}+|g\rangle_{SR}|0\rangle_{B_{1}}|0\rangle_{B_{2}}.$ (32)
Next, we apply $W^{(3)}_{B_{1}B_{2}}$, defined in Eq.(14), to $|\Phi_{3}\rangle$:
$|\Phi_{4}\rangle=W^{(3)}_{B_{1}B_{2}}|\Phi_{3}\rangle=\frac{1}{2^{n^{(R)}/2}}\sum_{i=0}^{N^{(S)}-1}\sum_{j=0}^{N^{(R)}-1}a_{ij}|i\rangle_{S}|0\rangle_{R}\Big(M_{B_{2}}|1\rangle_{B_{1}}|1\rangle_{B_{2}}\Big)+|g\rangle_{SR}|0\rangle_{B_{1}}|0\rangle_{B_{2}}.$ (33)
Using Eq.(22) with $\alpha\neq 0$, we stay with the first term involving $|1\rangle_{B_{2}}$ in (33) thus removing the garbage and resulting in
$|\Phi_{4}\rangle=|\Psi_{out}\rangle_{S}|0\rangle_{R}|1\rangle_{B_{1}}|1\rangle_{B_{2}},$ (34)
where $|\Psi_{out}\rangle_{S}$ and the normalization constant $G$ are defined in Eq.(7). This modification of the algorithm allows us to avoid the problem of small success probability for the ancilla-state measurement. It is remarkable that we do not even need to read the result of measurement because it is known in advance and, moreover, we do not use this result of measurement. This operation is involved just to remove the garbage from the superposition quantum state. The depth of the described algorithm is determined by the operator $W^{(1)}_{RB_{1}}$ and equals $O(n^{(R)})$. The circuit realizing this algorithm is given in Fig.1.
Figure 1: Application of controlled measurement in the algorithm present in Sec.II.1.
III.2 Controlled measurement in algorithms present in Refs.ZQKW_2024 ; ZBQKW_arXive2024
All the states obtained in Refs.ZQKW_2024 ; ZBQKW_arXive2024 before applying the final ancilla-state measurement have the form
$|\Phi\rangle=\alpha|E\rangle|1\rangle_{B_{1}}+\beta|F\rangle|0\rangle_{B_{1}},\;\;|\alpha|^{2}+|\beta|^{2}=1,\;\;|\alpha|\ll|\beta|,\;\;\langle F|E\rangle=0,\;\;\langle E|E\rangle=\langle F|F\rangle=1.$ (35)
where the state $|E\rangle$ contains the useful information, while the state $|F\rangle$ gathers all the garbage. In Eq.(35), the useful information and garbage are labeled by the states, respectively, $|1\rangle_{B_{1}}$ and $|0\rangle_{B_{1}}$ of the one-qubit ancilla $B_{1}$. To arrange the controlled measurement we, first of all, introduce the additional one-qubit ancilla $B_{2}$ in the ground state and apply the operator $W^{(2)}_{B_{1}B_{2}}$, defined in Eq.(12), to the state $|\Phi\rangle|0\rangle_{B_{2}}$:
$|\Phi_{1}\rangle=W^{(2)}_{B_{1}B_{2}}|\Phi\rangle|0\rangle_{B_{2}}=\alpha|E\rangle|1\rangle_{B_{1}}|1\rangle_{B_{2}}+\beta|F\rangle|0\rangle_{B_{1}}|0\rangle_{B_{2}}.$ (36)
After that we apply the controlled measurement $W^{(3)}_{B_{1}B_{2}}$, defined in Eq.(14), to $|\Phi_{1}\rangle$ thus selecting the term with the state $|1\rangle_{B_{2}}$ from the superposition present in Eq.(36) and obtain the following normalized resulting state, according to Eq.(22):
$|\Phi_{2}\rangle=W^{(3)}_{B_{1}B_{2}}|\Phi_{1}\rangle=\alpha|E\rangle\Big(M_{B_{2}}|1\rangle_{B_{1}}|1\rangle_{B_{2}}\Big)+\beta|F\rangle|0\rangle_{B_{1}}|0\rangle_{B_{2}}=\frac{\alpha}{|\alpha|}|E\rangle|1\rangle_{B_{1}}|1\rangle_{B_{2}}.$ (37)
In this way, the problem of small success probability in the matrix-manipulation algorithms can be removed. Scheme for replacing the usual measurement by the controlled one is shown in Fig.2.
Figure 2: Replacement of ordinary measurement of ancilla state (left circuit) with the subroutine of controlled measurement (right circuit).
III.3 Multiple random choice in superposition state
We consider the superposition state
$|\Phi_{0}\rangle=\sum_{j=1}^{Q}\alpha_{j}|E_{j}\rangle+\beta|F\rangle,\;\;\langle E_{j}|F\rangle=0,\;\;\langle E_{j}|E_{k}\rangle=\delta_{jk},\;\;\sum_{j=1}^{Q}|\alpha_{j}|^{2}+|\beta|^{2}=1,$ (38)
where $|E_{j}\rangle$, $j=1,\dots,Q$, and $|F\rangle$ are orthonormal states of some quantum system, and the states $|E_{j}\rangle$ contain the useful information. We label the states $|E_{j}\rangle$ and $|F\rangle$ using $(\lceil\log(Q+1)\rceil+1)$-qubit ancilla $B$:
$|\Phi_{0}\rangle\to|\Phi_{1}\rangle=\sum_{j=1}^{Q}\alpha_{j}|E_{j}\rangle|1\rangle_{B_{1}}|j\rangle_{B^{\prime}}+\beta|F\rangle|0\rangle_{B_{1}}|0\rangle_{B^{\prime}},$ (39)
where $B_{1}$ is the first cubit of the ancilla $B$ and $B^{\prime}$ is the ancilla $B$ without the first qubit. We assume that such labeling can be done using the particular structures of the states $|E_{j}\rangle$, see, for instance, Refs. ZQKW_2024 ; ZBQKW_arXive2024 .
Now we introduce the controlled measurement, similar to that given in (14),
$W_{B}=|1\rangle_{B_{1}}\,{{}_{B_{1}}\langle 1|}\otimes M_{B^{\prime}}+|0\rangle_{B_{1}}\,{{}_{B_{1}}\langle 0|}\otimes I_{B^{\prime}},$ (40)
where $M_{B^{\prime}}$ is the measurement operator applied to each qubit of $B^{\prime}$. Applying the controlled measurement $W_{B}$ to $|\Phi_{1}\rangle$ we fix $B^{\prime}$ in one of the states $|j\rangle_{B^{\prime}}$, $j\neq 0$, with the appropriate probability $p_{j}$, and this state is attached to the state $|E_{j}\rangle$. Thus, the garbage (the second term in Eq.(39)) disappears from the result, which is reflected in the following formula:
$|\Phi_{2}\rangle=W_{B}|\Phi_{1}\rangle=M_{B^{\prime}}\sum_{j=1}^{Q}\alpha_{j}|E_{j}\rangle|1\rangle_{B_{1}}|j\rangle_{B^{\prime}}+\beta|F\rangle|0\rangle_{B_{1}}|0\rangle_{B^{\prime}}={\sum_{j=1}^{Q}p_{j}|E_{j}\rangle|1\rangle_{B_{1}}|j\rangle_{B^{\prime}}.}$ (41)
In other words, the result of measurement $M_{B^{\prime}}$ yields the state $|E_{j}\rangle|1\rangle_{B_{1}}$ with the probability $p_{j}=|\alpha_{j}|^{2}/\sum_{j=1}^{Q}|\alpha_{j}|^{2}$, which does not depend on the probability amplitude $\beta$.
IV Extensions of matrix encoding technique
Both extensions of matrix-manipulation algorithms represented in Secs.IV.1 and IV.2 are based on the modification of matrix encoding into the initial state of a quantum system. Therefore, we first recall the matrix encoding into the initial state of the matrix-multiplication algorithm ZQKW_2024 ; ZBQKW_arXive2024 (the initial sate for the matrix-inversion algorithm ZBQKW_arXive2024 can be modified similarly).
The matrix encoding implemented in the algorithm for matrix multiplication uses two $n$-qubit quantum subsystems $R$ and $C$ whose basis states label, respectively, the rows and columns associated with each element $a_{jk}$ (complex in general) of the $N\times N$ matrix $A=\{a_{jk}:j,k=0,\dots,N-1\}$, $N=2^{n}$. Thus, the matrix $A$ is encoded as follows:
$|\Psi_{A}\rangle$ $=$ $\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}a_{jk}|j\rangle_{R}|k\rangle_{C}$ (42)
with the following normalization condition for the pure state of a quantum system:
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}=1.$ (43)
We introduce two extensions to the algorithms of matrix manipulations.The first of them is aimed on effective implementing the operations with complex matrices, Sec.IV.1. The second one allows to weaken the constraint, Eq.(43), imposed on the matrix elements by the normalization condition holding for pure quantum states, Sec.IV.2. We deal with the square $N\times N$ matrices for convenience. If the original matrix is rectangular we turn it to the square matrix by appending the appropriate number of either zero rows or zero columns. In Sec. IV.3, both above extensions are combined in the single circuit. Then, in Sec.V, both these extensions will be implemented in the matrix-multiplication algorithm. Therein, we also compare the results of application of the usual and controlled measurements.
IV.1 Hermitian conjugate
The complex elements of the $N\times N$ matrix $A=\{a_{jk}:j,k=0,\dots N-1\}$ can be written as
$a_{jk}=a_{jk0}+ia_{jk1}.$ (44)
We again involve two mentioned above $n$-qubit subsystems $R$ nd $C$ and add the third one-qubit subsystem $M$ whose state is $|0\rangle_{M}$ for the real part of $a_{jk}$ and $|1\rangle_{M}$ for the imaginary part. Then the elements $a_{jk}$ of the matrix $A$ can be encoded into the superposition state of three subsystems $R$, $C$ and $M$ as follows:
$|\Psi_{A}\rangle=\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}\sum_{m=0}^{1}a_{jkm}|j\rangle_{R}|k\rangle_{C}|m\rangle_{M}.$ (45)
With such pure state, the Hermitian conjugation of $A$ can be realized by the following operator:
$W=SWAP_{RC}\sigma^{(z)}_{M}$, (46)
where $SWAP_{RC}$ exchanges states of the subsystems $R$ and $C$ (in other words, it performs transposition of $A$), and $\sigma^{(z)}_{M}$ is the $\sigma^{(z)}$-operator applied to $M$. This operator changes the sign of the excited state $|1\rangle_{M}$ and thus performs the complex conjugation of $A$. Therefore, the matrix encoded into the state $W|\Psi_{A}\rangle$ is the Hermitian conjugate of $A$,
$|\tilde{\Psi}_{A}\rangle=W|\Psi_{A}\rangle=\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}\sum_{m=0}^{1}a_{kjm}(-1)^{m}|j\rangle_{R}|k\rangle_{C}|m\rangle_{M}.$ (47)
The Hermitian conjugation algorithm can be used as a subroutine in other algorithms, as shown in Sec.V.1. The circuit for the Hermitian conjugation is presented in Fig.3.
Figure 3: The circuit for the Hermitian conjugation, $Z\equiv\sigma^{(z)}$.
We shall notice that using the new encoding of the complex matrix in the superposition state of the triple system $R\otimes C\otimes M$ requires appropriate modifications for the multiplication algorithm which is considered in Sec.V. In addition, since the algorithms for determinant calculation and matrix inversion ZBQKW_arXive2024 use the row-wise matrix encoding, each row requires its own additional one-qubit subsystem to label the real and imaginary parts of its matrix elements, i.e., $N$ one-qubit subsystems $M_{i}$, $i=0,\dots,N-1$, are required for those algorithms.
IV.2 Weakened normalization constraint for matrix encoding
Eq.(43) represents a constraint imposed on the modulus of the matrix elements. However, we can weaken this constraint reducing it to the inequality
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}\leq 1.$ (48)
This can be done if we introduce, along with the subsystems $R$ and $C$ enumerating rows and columns of $A$, the one-qubit subsystem $K$ and replace the initial state given in Eq.(42) with the following one:
$|\Psi_{A}\rangle$ $=$ $b|0\rangle_{R}|0\rangle_{C}|0\rangle_{K}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}a_{jk}|j\rangle_{R}|k\rangle_{C}|1\rangle_{K}$ (49)
with the normalization
$|b|^{2}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}=1.$ (50)
Eq.(50) yields inequality (48). Thus, replacing Eq.(43) with inequality (48) becomes possible due to introducing the additional probability amplitude $b$ into the initial state, Eq.(49). Again, to weaken the normalization constraint in the row-wise encoding ZBQKW_arXive2024 , an additional qubit is requored for encoding each row.
IV.3 Circuit for new encoding of initial state
Now we combine the modifications of the state encoding presented in Eqs.(45) and (49) and encode the elements $a_{ij}$ decomposed into the real and imaginary parts, shown in Eq.(44), as follows:
$|\Psi_{A}\rangle$ $=$ $(b_{0}|0\rangle_{M}+b_{1}|1\rangle_{M})|0\rangle_{R}|0\rangle_{C}|0\rangle_{K}+$ (52)
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}(a_{jk0}|0\rangle_{M}+a_{jk1}|1\rangle_{M})|j\rangle_{R}|k\rangle_{C}|1\rangle_{K},\;\;b=b_{0}+ib_{1},$
$|b|^{2}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}=1.$
Replacement of matrix encoding via Eq. (42) with encoding via Eq.(52) is shown in Fig.4.
Figure 4: Replacement of matrix encoding via Eq.(42) with encoding via Eq.(52) for organizing effective manipulations with complex matrices and weakening the constraint on the matrix elements.
V Matrix multiplication
In this section we propose the matrix multiplication algorithm ZQKW_2024 ; ZBQKW_arXive2024 that takes into account the extensions discussed in Secs. IV, as well as implements the controlled measurement described in Sec.II.2.
Proposed modification is based on passing from encoding via Eq.(42) (used in ZQKW_2024 ) to encoding via Eq.(52) for the square $N\times N$ matrices $A^{(1)}$ and $A^{(2)}$ ($A^{(l)}=\{a^{(l)}_{jk}:j,k=0,\dots,N-1\}$, $l=1,2$) according to Sec.IV.
[…]
$|\Psi^{(l)}\rangle$ $=$ $(b^{(l)}_{0}|0\rangle_{M^{(l)}}+b^{(l)}_{1}|1\rangle_{M^{(l)}})|0\rangle_{R^{(l)}}|0\rangle_{C^{(l)}}|0\rangle_{K^{(l)}}+$ (54)
$\sum_{j,k=0}^{N-1}(a^{(l)}_{jk0}|0\rangle_{M^{(l)}}+a^{(l)}_{jk1}|1\rangle_{M^{(l)}})|j\rangle_{R^{(l)}}|k\rangle_{C^{(l)}}|1\rangle_{K^{(l)}},$
$|b^{(l)}|^{2}+\sum_{j,k}|a^{(l)}_{jk}|^{2}=1,\;\;l=1,2,$
where the constant parameters $b^{(l)}$ are introduced in accordance with Sec.IV.2. The initial state of the whole system reads
$|\Phi_{0}\rangle=|\Psi^{(2)}\rangle\otimes|\Psi^{(1)}\rangle=$ (55)
$\sum_{m_{1},m_{2}=0}^{1}b^{(1)}_{m_{1}}b^{(2)}_{m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}+$
$\sum_{{j_{1},k_{1},}\atop{j_{2},k_{2}=0}}^{N-1}\sum_{m_{1},m_{2}=0}^{1}\left(a^{(1)}_{j_{1}k_{1}m_{1}}a^{(2)}_{j_{2}k_{2}m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}\right)$
$\times|j_{1}\rangle_{R_{1}}|k_{1}\rangle_{C^{(1)}}|j_{2}\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{0}\rangle,$
where we select the terms with the states $|j\rangle_{K^{(1)}}|j\rangle_{K^{(2)}}$, $j=0,1$, which are needed below, while the terms with the states $|j\rangle_{K^{(1)}}|k\rangle_{K^{(2)}}$, $j\neq k$, are collected in the garbage $|g_{0}\rangle$.
From the superposition state $|\Phi_{0}\rangle$, we have to select the terms with the states $|k\rangle_{C^{(1)}}|k\rangle_{R^{(2)}}$, $k=0,\dots,N-1$, i.e., $k_{1}=j_{2}$ in the second part in the right-hand side of Eq.(55). For that purpose we apply the CNOTs $W_{j}$ to the $j$th qubits of $C^{(1)}$ and $R^{(2)}$,
$W_{j}$ $=$ $|1\rangle_{C^{(1)}_{j}}\;{}_{C^{(1)}_{j}}\langle 1|\otimes\sigma^{(x)}_{R^{(2)}_{j}}+|0\rangle_{C^{(1)}_{j}}\;{}_{C^{(1)}_{j}}\langle 0|\otimes I_{R^{(2)}_{j}},\;\;j=1,\dots,n.$ (56)
where $\sigma^{(x)}_{R^{(2)}_{j}}$ and $I_{R^{(2)}_{j}}$ are, respectively, the $\sigma^{(x)}$-operator and identity operator applied to the $j$th qubit of the subsystem $R^{(2)}$. We have chosen the qubits of $C^{(1)}$ as controlling ones. All operators $W_{j}$, $j=1,\dots,n$, commute with each other. Applying the operator $W^{(0)}_{C^{(1)}R^{(2)}}$,
$W^{(0)}_{C^{(1)}R^{(2)}}=\prod_{j=1}^{n}W_{j},$ (57)
to $|\Phi_{0}\rangle$ we obtain:
$|\Phi_{1}\rangle=W^{(0)}_{C^{(1)}R^{(2)}}|\Phi_{0}\rangle=$
$\sum_{m_{1},m_{2}=0}^{1}b^{(1)}_{m_{1}}b^{(2)}_{m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}+$
$\sum_{{j_{1},j,}\atop{k_{2}=0}}^{N-1}\sum_{m_{1},m_{2}=0}^{1}\left(a^{(1)}_{j_{1}jm_{1}}a^{(2)}_{jk_{2}m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}\right)$
$\times|j_{1}\rangle_{R^{(1)}}|j\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{1}\rangle.$
The first and second parts in the right-hand side of Eq.(V) collect the terms with the state $|0\rangle_{R_{2}}|j\rangle_{K^{(1)}}|j\rangle_{K^{(2)}}$, $j=0,1$, (the other terms are garbage by definition, they are collected in $|g_{1}\rangle$ above). We note that the first part in the right-hand side of Eq.(V) remains unchanged after applying the operator $W^{(0)}_{C^{(1)}R^{(2)}}$. We also note that the operator $W^{(0)}_{C^{(1)}R^{(2)}}$ transforms $|g_{0}\rangle$ as well. However, we don’t need details of that transformation. The similar statement can be referred to the operators $W^{(1)}_{C^{(1)}}$, $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ and $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B}$ below.
Next, we apply the Hadamard transformations
$W^{(1)}_{C^{(1)}}=H^{\otimes n}$ (59)
to the qubits of $C^{(1)}$ and select the terms with $|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|j\rangle_{K^{(1)}}|j\rangle_{K^{(2)}}$, $j=0,1$, putting others to the garbage $|g_{2}\rangle$:
$|\Phi_{2}\rangle=W^{(1)}_{C^{(1)}}|\Phi_{1}\rangle=$ (60)
$\frac{1}{2^{n/2}}\sum_{m_{1},m_{2}=0}^{1}b^{(1)}_{m_{1}}b^{(2)}_{m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}$
$+\frac{1}{2^{n/2}}\sum_{j_{1},k_{2}=0}^{N-1}\sum_{j=0}^{N-1}\sum_{m_{1},m_{2}=0}^{1}\left(a^{(1)}_{j_{1}jm_{1}}a^{(2)}_{jk_{2}m_{2}}|m_{1}\rangle_{M^{(1)}}|m_{2}\rangle_{M^{(2)}}\right)$
$\times|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{2}\rangle.$
Now, to complete multiplication, we have to organize proper combinations of products of real and imaginary parts of matrices $A^{(i)}$, $i=1,2$, to form the real and imaginary parts of the result. To this end we introduce the following controlled operator:
$W_{M^{(1)}M^{(2)}}=|1\rangle_{M^{(2)}}\,{}_{M^{(2)}}\langle 1|\otimes\sigma^{(x)}_{M^{(1)}}\sigma^{(z)}_{M^{(1)}}+|0\rangle_{M^{(2)}}\,{}_{M^{(2)}}\langle 0|\otimes I_{M^{(1)}}.$ (61)
We also introduce the controlled operator
$W_{K^{(1)}K^{(2)}}=|1\rangle_{K^{(1)}}\,{}_{K^{(1)}}\langle 1|\otimes\sigma^{(x)}_{K^{(2)}}+|0\rangle_{K^{(1)}}\,{}_{K^{(1)}}\langle 0|\otimes I_{K^{(2)}}$ (62)
which labels all needed terms by the state $|0\rangle_{K^{(2)}}$. We include the operators $W_{M^{(1)}M^{(2)}}$ and $W_{K^{(1)}K^{(2)}}$ into the operator
$W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}=W_{K^{(1)}K^{(2)}}H_{M^{(2)}}W_{M^{(1)}M^{(2)}},$ (63)
where $H_{M^{(2)}}$ is the Hadamard operator applied to the single qubit of $M^{(2)}$, and apply $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ to $|\Phi_{2}\rangle$, thus obtaining
$|\Phi_{3}\rangle=W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}|\Phi_{2}\rangle=$ (64)
$\frac{1}{2^{(n+1)/2}}\Big((b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1})|0\rangle_{M^{(1)}}+(b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0})|1\rangle_{M^{(1)}}\Big)$
$\times|0\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}$
$+\frac{1}{2^{(n+1)/2}}\sum_{j_{1},k_{2}=0}^{N-1}\sum_{j=0}^{N-1}\left((a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}0}-a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}1})|0\rangle_{M^{(1)}}\right$.
$\left.+(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}1}+a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}0})|1\rangle_{M^{(1)}}\right)$
$\times|0\rangle_{M^{(2)}}|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}+|g_{3}\rangle.$
Here the first term plays an auxiliary role similar to the first term of Eq.(54), while the second term contains the desired matrix product. All extra terms are collected in the garbage $|g_{3}\rangle$. We emphasize that the operator $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ is a novelty in comparison with multiplication algorithm in ZBQKW_arXive2024 .
Next, to label and remove the garbage, we prepare the two-qubit ancillae $B$ in the ground state $|0\rangle_{B}$, introduce the projector
$P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}=|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{M^{(2)}}|0\rangle_{K^{(2)}}\;{}_{C^{(1)}}\langle 0|_{R^{(2)}}\langle 0|_{M^{(2)}}\langle 0|_{K^{(2)}}\langle 0|$ (65)
and the controlled operator $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B}$,
$W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B}=P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}\otimes\sigma^{(x)}_{B_{1}}\sigma^{(x)}_{B_{2}}$ (66)
$+(I_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}-P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}})\otimes I_{B},$
of the depth $O(n)$ with $2(n+1)$-qubit control register. Here $B_{i}$, $i=1,2$, means the $i$th qubit of the ancilla $B$. Applying this operator to $|\Phi_{3}\rangle|0\rangle_{B}$ we obtain
$|\Phi_{4}\rangle=W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B}|\Phi_{3}\rangle|0\rangle_{B}=$ (67)
$\frac{1}{2^{(n+1)/2}}\Big((b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1})|0\rangle_{M^{(1)}}+(b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0})|1\rangle_{M^{(1)}}\Big)$
$\times|0\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}|1\rangle_{B_{1}}|1\rangle_{B_{2}}+$
$\frac{1}{2^{(n+1)/2}}\sum_{{j_{1},k_{2},}\atop{j=0}}^{N-1}\Big((a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}0}-a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}1})|0\rangle_{M^{(1)}}+(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}1}+a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}0})|1\rangle_{M^{(1)}}\Big)$
$\times|0\rangle_{M^{(2)}}|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}\ |1\rangle_{B_{1}}|1\rangle_{B_{2}}+|g_{3}\rangle|0\rangle_{B_{1}}|0\rangle_{B_{2}}.$
Now we have to remove the garbage marked by the state $|0\rangle_{B_{1}}|0\rangle_{B_{2}}$. Before implementing the controlled measurement described in Sec.II.2.2, we apply the most straightforward way of garbage removal that is the usual measurement of the state of $B_{2}$ with the desired output $|1\rangle_{B_{2}}$. Of course, this way requires multiple running of the algorithm because the single run with the subsequent measurement might yield undesired result $|0\rangle_{B_{2}}$ that must be disregarded. But once the desired output is obtained, the quantum state of the remaining system becomes
$|\Phi_{5}\rangle=|\Psi_{out}\rangle\,|0\rangle_{M_{2}}|0\rangle_{C_{1}}|0\rangle_{R_{2}}|0\rangle_{K_{2}}|1\rangle_{B_{1}},$ (68)
$|\Psi_{out}\rangle=G^{{-1}}\Big(\Big((b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1})|0\rangle_{M^{(1)}}+(b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0})|1\rangle_{M^{(1)}}\Big)$ (69)
$\times|0\rangle_{R^{(1)}}|0\rangle_{C^{(2)}}|0\rangle_{K_{1}}+$
$\sum_{{j_{1},k_{2},}\atop{j=0}}^{N-1}\left((a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}0}-a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}1})|0\rangle_{M^{(1)}}+(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}1}+a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}0})|1\rangle_{M^{(1)}}\right)$
$\times|j_{1}\rangle_{R^{(1)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}\Big)=$
$G^{-1}\left((\hat{b}_{0}|0\rangle_{M^{(1)}}+\hat{b}_{1}|1\rangle_{M^{(1)}})|0\rangle_{R^{(1)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}\right$.
$\left.+\sum_{j,k=0}^{N-1}(\hat{a}_{jk0}|0\rangle_{M^{(1)}}+\hat{a}_{jk1}|1\rangle_{M^{(1)}})|j\rangle_{R^{(1)}}|k\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}\right),$
where $G=\left(|b^{(1)}b^{(2)}|^{2}+\sum_{j,k=0}^{N-1}\Big|\sum_{l=0}^{N-1}a^{(1)}_{jl}a^{(2)}_{lk}\Big|^{2}\right)^{1/2}$ is the normalization factor,
$\hat{b}_{0}=b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1},\;\;\hat{b}_{1}=b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0}$, (70)
$\hat{a}_{jk0}=\sum_{l=0}^{N-1}(a^{(1)}_{jl0}a^{(2)}_{lk0}-a^{(1)}_{jl1}a^{(2)}_{lk1}),\;\;\hat{a}_{jk1}=\sum_{l=0}^{N-1}(a^{(1)}_{jl0}a^{(2)}_{lk1}+a^{(1)}_{jl1}a^{(2)}_{lk0})$,
so that $\hat{A}=A^{(1)}A^{(2)}=\{\hat{a}_{jk}:j,k=0,\dots,N-1\}$, $\hat{a}_{jk}=\hat{a}_{jk0}+i\hat{a}_{jk1}$, $\hat{b}=\hat{b}_{0}+i\hat{b}_{1}$, where $\hat{a}_{jkm}$, $\hat{b}_{m}$, $m=0,1$, are real numbers. We note that the first term in the state $|\Psi_{out}\rangle$ given in Eq.(69) plays the same role as the first term in the expressions for $|\Psi^{(l)}\rangle$, given in Eq.(54). It weakens the constraint on the normalization of the elements of the resulting matrix $\hat{A}$.
The weak point is that the probability of measuring the desired state $|1\rangle_{B_{2}}$ is small, $p=G^{2}/2^{n+1}$, and exponentially decreases with an increase in $n$. Therefore, to obtain the desired result $|1\rangle_{B_{2}}$ we have to run the algorithm $O(2^{n+1}/G^{2})$ times. However, we can replace the usual measurement by the controlled one, i.e., let us perform the controlled measurement of the ancilla qubit $B_{2}$ via the operator (see also Eq.(14) in Sec.II.2.2)
$W^{(4)}_{B}=|1\rangle_{B_{1}}\,{}_{B_{1}}\langle 1|\otimes M_{B_{2}}+|0\rangle_{B_{1}}\,{}_{B_{1}}\langle 0|\otimes I_{B_{2}},$ (71)
where $M_{B_{2}}$ is the measurement operator applied to the qubit $B_{2}$. Then, applying $W^{(4)}_{B}$ to $|\Phi_{4}\rangle$ , we select the term with the state $|1\rangle_{B_{2}}$ from this superposition, thus removing the garbage and obtaining:
$|\Phi_{5}\rangle=W^{(4)}_{B}|\Phi_{4}\rangle=|\Psi_{out}\rangle\,|0\rangle_{M_{2}}|0\rangle_{C_{1}}|0\rangle_{R_{2}}|0\rangle_{K_{2}}|1\rangle_{B_{1}}|1\rangle_{B_{2}},$ (72)
where $|\Psi_{out}\rangle$ is determined in Eq.(69). In this case, only the single run of the algorithm is enough to get the desired state $|\Psi_{out}\rangle$.
Figure 5: (a) The circuit for the matrix multiplication algorithm; the state $|\Psi_{out}\rangle$ is formed by the subsystems $R_{1}$, $C_{2}$, $K_{1}$ and $M_{1}$; $X\equiv\sigma^{(x)}$, $Z\equiv\sigma^{(z)}$; we omit subscripts in notations $W^{(j)}$ for the brevity; operator $W^{(2)}$ is a novelty in comparison with earlier algorithm in ZBQKW_arXive2024 ; (b) Notation for the multi-qubit CNOT.
V.1 Manipulations with input matrices
In this section we consider several manipulations with the input matrices $A^{(1)}$ and $A^{(2)}$. Instead of multiplication $A^{(1)}A^{(2)}$ we can perform multiplications $(A^{(1)})^{\dagger}A^{(2)}$, $A^{(1)}(A^{(2)})^{\dagger}$, $(A^{(1)})^{\dagger}(A^{(2)})^{\dagger}$, $(A^{(1)}A^{(2)})^{\dagger}=(A^{(2)})^{\dagger}(A^{(1)})^{\dagger}$ and change the order of the multipliers. With this purpose we introduce the following operators.
To organize the Hermitian conjugate of $A^{(i)}$, $i=1,2$, we use operator $W$, Eq.(46), introduced in Sec.IV.1 and rewrite it using the new notations:
$Q^{(i)}_{R^{(i)}C^{(i)}M^{(i)}}=SWAP_{R^{(i)}C^{(i)}}\sigma^{(z)}_{M^{(i)}},\;\;i=1,2.$ (73)
To change the order of the matrices in the multiplication algorithm we apply the operator
$Q^{(3)}_{R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}K^{(1)}K^{(2)}}=SWAP_{R^{(1)}R^{(2)}}SWAP_{C^{(1)}C^{(2)}}SWAP_{M^{(1)}M^{(2)}}SWAP_{K^{(1)}K^{(2)}}.$ (74)
Here, the multi-qubit $SWAP$-operators $SWAP_{R^{(1)}R^{(2)}}$ and $SWAP_{C^{(1)}C^{(2)}}$ establish exchange between the appropriate qubits of two selected subsystems. The circuit implementing these operators is shown in Fig.6. For instance, to calculate the product $A^{(1)}(A^{(2)})^{\dagger}$ we apply the operator $Q^{(2)}_{R^{(2)}C^{(2)}M^{(2)}}$ to the state $|\Phi_{0}\rangle$, Eq.(55). To calculate $(A^{(1)}A^{(2)})^{\dagger}$ we apply the operators $Q^{(3)}_{R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$, $Q^{(2)}_{R^{(2)}C^{(2)}M^{(2)}}$, $Q^{(1)}_{R^{(1)}C^{(1)}M^{(1)}}$ to the state $|\Phi_{0}\rangle$. Application of operators $Q^{(i)}$, $i=1,2,3$, can be controlled by additional qubits $q_{i}$ of the subsystem $q$ (see Fig.6), i.e., by the following controlled operators:
$\hat{Q}^{(i)}_{S^{(i)}}=|1\rangle_{q_{i}}\,{}_{q_{i}}\langle 1|\otimes Q^{(i)}_{S^{(i)}}+|0\rangle_{q_{i}}\,{}_{q_{i}}\langle 0|\otimes I_{S^{(i)}},$ (75)
where $S^{(i)}=R^{(i)}C^{(i)}M^{(i)}$, $i=1,2$, $S^{(3)}=R^{(1)}C^{(1)}R^{(2)}C^{(2)}M^{(1)}M^{(2)}K^{(1)}K^{(2)}$ and $I_{S^{(i)}}$ is the identity operator applied to $S^{(i)}$. The state $|\varphi\rangle_{q}=|n_{1}\rangle_{q_{1}}|n_{2}\rangle_{q_{2}}|n_{3}\rangle_{q_{3}}$, $n_{i}=0,1$ ($i=1,2,3$), of the controlling subsystem $q$ is not changed after running the algorithm.
Of course calculating $(A^{(1)}A^{(2)})^{\dagger}$ can be done by performing the Hermitian conjugation after multiplication $A^{(1)}A^{(2)}$, this is another option.
Figure 6: (a) Controlled manipulations with input matrices $A^{(1)}$ and $A^{(2)}$; (b) Notation for the multi-qubit SWAP-operator
V.2 Measuring normalization $G$ and discussion
Although the controlled measurement allows to resolve the problem of exponential decrease of the success probability with the number of qubits encoding considered matrices, we lose the information about normalization $G$ in Eq.(69). However, this normalization might be required in obtaining the final result. Below we explore the method of probabilistic measuring the normalization $G$.
Let $S_{1}=|b^{(1)}b^{(2)}|^{2}$, this value is known due to the normalization in matrix encoding, see Eq.(50). If, along with $S_{1}$, we know $S_{1}/G^{2}=\tilde{S}_{1}$, then $G=\sqrt{S_{1}/\tilde{S}_{1}}$. To find $\tilde{S}_{1}$, we refer to the state $|\Psi_{out}\rangle$, Eq.(69), and measure the state of the subsystem $K_{1}$ with the output $|0\rangle_{K_{1}}$. It is easy to understand that the probability of this output equals $\tilde{S}_{1}$, and can be found in result of multiple runs of the algorithm. It is important that the problem of small success probability is resolved on passing from $|\Phi_{4}\rangle$, Eq.(67), to $|\Phi_{5}\rangle$, Eq.(68), and does not appear in the above measurement, where the success probability $\tilde{S}_{1}$ does not depend on $n$.
We have to note that the last measurement destroys the part of superposition quantum state in which the result of matrix multiplication is stored. Therefore, to get the full information, i.e., the normalization $G$ and matrix product, we have to use the set of runs to probabilistically determine $G$ and then run the algorithm one more time to end up with the state $|\Phi_{5}\rangle$ that includes $|\Psi_{out}\rangle$ encoding the product $A^{(1)}A^{(2)}$. However, if $G$ is not needed, then the single run is enough.
We emphasize that both introduced extensions do not decline the parameters of the multiplication circuit such as depth and space. Fig.5 shows that the depth is determined by the operator $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}$, while the manipulations with the input matrices in Sec.V.1, see Fig.6, increase this depth only by insignificant constant factor. Thus, the estimation of the depth remains $O(n)$ (where $N=2^{n}$) similar to the estimation in ZBQKW_arXive2024 . Regarding the space, the circuit in Fig.5 includes only five additional one-qubit subsystems ($K^{(i)}$, $M^{(i)}$, $i=1,2$ , and $B_{2}$) in comparison with the circuit in ZBQKW_arXive2024 with one of them (the ancilla qubit $B_{2}$) is needed for organization of controlled measurement. Therefore, the space estimation remains $O(n)$ as well. Of course, multiple running is required in the case of implementing the usual ancilla-measurement ZBQKW_arXive2024 . Then the depth of the whole algorithm is bigger than the depth of the circuit by $O(2^{n+1}/G^{2})$ times and equals $O(2^{n+1}n/G^{2})\sim O(2^{n}n)$.
VI Conclusion
In this paper we study some new aspects of quantum algorithms with their application to the matrix-manipulation algorithms proposed in ZQKW_2024 ; ZBQKW_arXive2024 and based on the encoding the matrix elements into the pure superposition state of a quantum system. However, applicability of introduced novelties spreads much beyond of proposed examples.
The most important issue is the replacement of the ordinary one-qubit ancilla-state measurement (needed to remove the garbage acquired during evaluation of the quantum algorithm) with the measurement of the ancilla-state controlled by the state of another one-qubit ancilla (controlled measurement). This modification allows to resolve the problem of small access probability to the desired state of the ancilla. Remember, that the problem of small success probability appears in the case of ordinary measurement of the ancilla state in HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 ; ZQKW_2024 ; ZBQKW_arXive2024 . Thus, the controlled measurement significantly increases the efficiency of the algorithms developed in those papers. At that, we lose the information on the normalization constant for the output state because this constant is expressed in terms of the access probability to the desired ancilla state and this probability is not measured in the modified algorithm. However, the normalization constant can be revealed as shown in the matrix multiplication algorithm, Sec.V.2. We also have to note that the method of practical realization of the controlled measurement is not developed yet, but its realization would be one more justification of reality of superposition quantum states.
We also introduce two extensions for the quantum data encoding and apply them to the matrix-multiplication algorithm proposed in ZBQKW_arXive2024 . The first extension consists in separating the real and imaginary parts of the complex matrix under consideration by encoding these two parts into two orthogonal subspaces of the quantum state-space. Such extension allows constructing effective subroutine for operations typical for complex numbers, which are complex and Hermitian conjugations. However, this separation of real and imaginary parts requires proper modifications of the quantum algorithms, in particular, quantum algorithms for matrix addition, multiplication, determinant calculation, matrix inversion and solving linear systems that are described in ZQKW_2024 ; ZBQKW_arXive2024 . In our paper, we present such modification for the multiplication algorithm and demonstrate manipulations with the input matrices using Hermitian conjugation subroutine.
Our second extension weakens the normalization constraint, imposed on the elements of the input matrices, which follows from the normalization condition for a superposition pure quantum state. This can be done by including an additional term in the superposition state encoding a particular matrix which leads to replacing the normalization equality, Eq.(43), with inequality (48) that imposes only the upper bound on the sum of the squared modulus of the matrix elements. We emphasize that the applicability of the two last extensions is not restricted to the matrix encoding, any data-set can be used instead of matrices.
Of course, each of the above modifications requires involving an extra qubit. Therefore, along with four subsystems $R^{(i)}$ and $C^{(i)}$, $i=1,2$, enumerating rows and columns of the matrices $A^{(i)}$, $i=1,2$, the $i$th matrix is associated with the one-qubit subsystem $M^{(i)}$, responsible for separate encoding of the real and imaginary parts of complex matrix elements, and the one-qubit subsystem $K^{(i)}$ allowing to introduce an extra term into the superposition state encoding the matrix $A^{(i)}$, namely the presence of this extra term weakens the normalization constraint. Thus, only four additional qubits appear in the modified multiplication algorithm. In addition, involving the controlled measurement requires an additional qubit for the ancilla $B$ which becomes the two-qubit ancilla.
We emphasize that all the above modifications do not significantly decline the depth and space of the circuit for multiplication algorithm. Estimations for both of them remain $O(n)$, which is similar to the appropriate characteristics of the circuit for the multiplication algorithm in ZBQKW_arXive2024 . However, the depth of the whole multiplication algorithm in ZBQKW_arXive2024 is much bigger because of the multiple running of the algorithm to access the desired ancilla-state, it equals $O(2^{n}n)$.
Proposed modifications can be included in other matrix-manipulation algorithms considered in Refs.ZQKW_2024 ; ZBQKW_arXive2024 , as well as in any measurement-based algorithm HHL ; WWLN ; ZDB ; ZBD ; Wei ; WeiBook and algorithm based on data encoding ZZRF ; BWPRWL ; A ; RML ; WBL ; SSP ; Wang2 .
Acknowledgments. This project is supported by the National Natural Science Foundation of China (Grants No. 12031004, No. 12271474 and No. 61877054). The work was also partially funded by a state task of Russian Fundamental Investigations (State Registration No. 124013000760-0).